### v2

# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

Abstract.
In this paper, we compare a certain field naturally arising from the kernel of the pro-$p$ outer Galois representation of a once-punctured elliptic curve with complex multiplication over an imaginary quadratic field $K$ with the maximal pro-$p$ Galois extension of the mod-$p$ ray class field $K(p)$ of $K$ unramified outside $p$. Among other things, we prove that the two fields coincide with each other for every split prime $p$ satisfying certain conditions, including an analogue of the so-called Deligne-Ihara conjecture. This result provides an analogue of Sharifi’s result on the kernel of the pro-$p$ outer Galois representation of the thrice-punctured projective line.
Keywords. once-punctured elliptic curve, complex multiplication, étale fundamental group, Galois representation.
Mathematics Subject Classification : 11G05, 11G15, 11R32, 12E30.
1. Introduction
In this paper, we study the kernel of the pro-$p$ outer Galois representations associated to once-punctured elliptic curves with complex multiplication.
We recall the definition of the pro-$p$ outer Galois representations: Suppose that $X$ is a geometrically connected algebraic variety defined over a number field $F$. Throughout this paper, we work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$, and every number field is considered as a subfield of $\bar{\mathbb{Q}}$.
We denote the étale fundamental group of $X$ with respect to a (possibly tangential) basepoint $\bar{x}$ by $\pi_{1}(X,\bar{x})$. In the following, we write $\bar{X}\coloneqq X\times_{F}\bar{\mathbb{Q}}$. There is an exact sequence determined by the structure morphism $X\to\mathrm{Spec}(F)$:
$1\to\pi_{1}(\bar{X},\bar{x})\to\pi_{1}(X,\bar{x})\to G_{F}\coloneqq\mathrm{Gal}(\bar{\mathbb{Q}}/F)\to 1$
Hence the exact sequence, together with the conjugation action of $\pi_{1}(X,\bar{x})$ on $\pi_{1}(\bar{X},\bar{x})$, determines the outer Galois representation
$\rho_{X}\colon G_{F}\to\mathrm{Out}(\pi_{1}(\bar{X},\bar{x}))\coloneqq\mathrm{Aut}(\pi_{1}(\bar{X},\bar{x}))/\mathrm{Inn}(\pi_{1}(\bar{X},\bar{x}))$
which does not depend on the choice of basepoints. For a prime $p$, since the maximal pro-$p$ quotient $\pi_{1}(\bar{X},\bar{x})^{(p)}$ of $\pi_{1}(\bar{X},\bar{x})$ is characteristic, the outer representation $\rho_{X}$ induces a homomorphism
$\rho_{X,p}\colon G_{F}\to\mathrm{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)}),$
which we call the pro-$p$ outer Galois representation associated to $X$.
If $X$ is a hyperbolic curve, such outer representations are mainly studied in the context of anabelian geometry. Among other things, it is known that $\rho_{X}$ is injective if $X$ is a hyperbolic curve, cf. Matsumoto [Mat96] when $X$ is affine and Hoshi-Mochizuki [HM11] when $X$ is proper.
As for the pro-$p$ outer Galois representation, it is far from being injective since the group $\mathrm{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)})$ contains an open pro-$p$ subgroup. In particular, the fixed field of the kernel of $\rho_{X,p}$ is an almost pro-$p$ extension over $F$, and it seems to be interesting question to study arithmetic properties of this field.
In the case of the thrice-punctured projective line $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\}$, Anderson and Ihara [AI88, Theorem 2 (IV)] proved that the fixed field of the kernel of the pro-$p$ outer Galois representation is a nonabelian infinite pro-$p$ extension over the field $\mathbb{Q}(\mu_{p^{\infty}})$ unramified outside $p$, and is generated by certain $p$-units, which they call higher circular $p$-units, constructed in a certain geometric way. Moreover, they posed the following question [AI88, page 272, (a)]:
Is ¯$\bar{\mathbb{Q}}^{\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p})}$ the maximal pro-$p$ extension of $\mathbb{Q}(\mu_{p})$ unramified outside $p$?
Let $\Omega^{\mathrm{cyc}}$ denote the maximal pro-$p$ extension of the $p$-th cyclotomic field $\mathbb{Q}(\mu_{p})$ unramified outside $p$. Regarding Anderson-Ihara’s question, Sharifi obtained the following affirmative answer for odd regular primes, assuming a certain conjecture which is nowadays referred to as the Deligne-Ihara conjecture:
Theorem (Sharifi [Sha02, Theorem 1.1]).
Assume that $p$ is an odd regular prime and the Deligne-Ihara conjecture holds for $p$. Then the fixed field ¯$\bar{\mathbb{Q}}^{\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p})}$ coincides with $\Omega^{\mathrm{cyc}}$.
The Deligne-Ihara conjecture concerns with the structure of a graded Lie algebra over $\mathbb{Q}_{p}$ associated to a certain descending central filtration on $G_{\mathbb{Q}}$ induced by the pro-$p$ outer Galois representation of $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\}$, and states that this graded Lie algebra is freely generated by Soulé elements $\{\sigma_{m}\}_{m\geq 3,\mathrm{odd}}$ in each odd degree $\geq 3$. We do not go into further details (see Sharifi [Sha02] and Ihara [Iha02] for detailed discussions). However, we will formulate an analogue of the Deligne-Ihara conjecture for once-punctured elliptic curves with complex multiplication (Conjecture 2.8).
The Deligne-Ihara conjecture is now a theorem: The generation portion of the conjecture is proved by Hain-Matsumoto [HM03], and later Brown [Bro12] prove the conjecture in full generality. Combining Sharifi’s work with the settlement of the Deligne-Ihara conjecture, Anderson-Ihara’s question is affirmative if $p$ is odd and regular.
Remark 1.1.
If $p$ is odd, the Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is generated by $(p+1)/2$ elements if and only if $p$ is regular by [NSW08, (10.7.13) Theorem]. This observation plays an essential role in the proof of [Sha02, Theorem 1.1]
If $p$ is odd and regular, the Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is even a free pro-$p$ group of rank $(p+1)/2$ by loc. cit. and we have an isomorphism
$\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p^{\infty}}))^{\mathrm{ab}}\cong\mathbb{Z}_{p}[[\mathrm{Gal}(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}(\mu_{p}))]]^{\oplus\frac{p-1}{2}}$
between $\mathbb{Z}_{p}[[\mathrm{Gal}(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}(\mu_{p}))]]$-modules by [NQD84, Corollaire 2.7], together with the fact that the left-hand side is generated by $(p-1)/2$ elements as an Iwasawa module.
Remark 1.2.
We remark that there exists a pair $(X,p)$ consisting of a hyperbolic curve $X$ over $\mathbb{Q}$ of positive genus and a prime $p$ such that the kernel of the pro-$p$ outer Galois representation $\rho_{X,p}$ coincides with that of the thrice-punctured projective line. For example, the Fermat curve $x^{p}+y^{p}+1=0$ of degree $p$ satisfies this condition [Hos15, Example 3.4 (i)]. In particular, for such a pair $(X,p)$, the fixed field of the kernel of $\rho_{X,p}$ coincides with $\Omega^{\mathrm{cyc}}$ if $p$ is odd and regular. In general, for every pair $(X,p)$ as above, it is known by Hoshi-Mochizuki [HM11, Theorem C (i)] that
$\ker(\rho_{X,p})\subset\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\},p}).$
The aim of this paper is to prove an analogue of Sharifi’s result in the case of once-punctured elliptic curves with complex multiplication.
Let $K$ be an imaginary quadratic field of class number one, and $(E,O)$ an elliptic curve over $K$ with its origin $O\in E(K)$ having complex multiplication by the ring of integers $O_{K}$ of $K$. We assume that the complex multiplication by $O_{K}$ is defined over $K$, i.e. $O_{K}\hookrightarrow\mathrm{End}_{K}(E)$. In the following, for a prime $p$, we denote the mod-$p$ and mod-$p^{\infty}$ ray class fields of $K$ by $K(p)$ and $K(p^{\infty})$, respectively. The extension $K(p^{\infty})/K(p)$ is a $\mathbb{Z}_{p}^{2}$-extension.
The following main result characterizes the kernel of the pro-$p$ outer Galois representation associated to the complement $X\coloneqq E-O$ of the origin of the elliptic curve, which we call the once-punctured elliptic curve.
Theorem A (Theorem 2.12).
Let $p\geq 5$ be a prime splitting into two distinct primes in $K$. Assume that
  1. (1)
    the class number of the mod-$p$ ray class field $K(p)$ is not divisible by $p$,
  2. (2)
    there are only two primes of the mod-$p^{\infty}$ ray class field $K(p^{\infty})$ lying above $p$, and
  3. (3)
    an analogue of the Deligne-Ihara conjecture (Conjecture 2.8) holds.
Then the fixed field ¯$\bar{\mathbb{Q}}^{\ker(\rho_{X,p})}$ of the kernel of the pro-$p$ outer Galois representation of $X=E-O$ is equal to the compositum of the field $K(E[p])$ generated by the coordinates of $p$-torsion points of $E$ and the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
To the author’s knowledge, the pair $(X,p)$ as above gives a first example where the fixed field is determined independently of $(\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p)$, although it essentially depends on Conjecture 2.8.
In the following, we denote the maximal pro-$p$ extension of $K(p)$ unramified outside $p$ by $\Omega$. Note that $\Omega$ contains $\Omega^{\mathrm{cyc}}$, since $K(p)$ contains $\mathbb{Q}(\mu_{p})$.
Remark 1.3.
Let $p\geq 5$ be a prime which splits in $K$.
  1. (1)
    By [NSW08, (10.7.13) Theorem], we have
    $\dim_{\mathbb{F}_{p}}H^{1}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p})=[K(p):K]+s_{p}+\dim_{\mathbb{F}_{p}}Cl_{p}(K(p))/p,$
    $\dim_{\mathbb{F}_{p}}H^{2}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p})=-1+s_{p}+\dim_{\mathbb{F}_{p}}Cl_{p}(K(p))/p.$
[…]
  2. (2)
    The first and second assumptions of Theorem A allow us to determine the $\Lambda$-module structure of the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$, where
    $\Lambda\coloneqq\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$
    is the completed group ring over $\mathbb{Z}_{p}$ which is (non-canonically) isomorphic to the two-variable power series ring over $\mathbb{Z}_{p}$. More precisely, under these assumptions, we have
    $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}\cong\Lambda^{[K(p):K]-1}\oplus\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1)),$
    where $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\subset\Lambda$ is the annihilator ideal of the Tate twist $\mathbb{Z}_{p}(1)$, cf. Remark 4.3.
  3. (3)
    We do not know whether there are infinitely many primes satisfying the first assumption of Theorem A, because this assumption for $p$ implies that $p$ is regular. We also do not know whether the second assumption is satisfied by infinitely many primes. For example, there are 39175 prime numbers which are less than $10^{6}$ and split in $\mathbb{Q}(\sqrt{-1})$. Among them, we have checked that 13705 primes satisfy the second assumption.
Example 1.1.
We have checked whether a small prime $p$ satisfies the first two assumptions of Theorem A, by using SageMath [The24]. For example, when $p=5$, both assumptions are satisfied if and only if $K=\mathbb{Q}(\sqrt{-1})$. Similarly, when $p=7$, both conditions are satisfied only for $K=\mathbb{Q}(\sqrt{-3})$ or $\mathbb{Q}(\sqrt{-19})$. When $p=13$ or $17$, we also checked that both assumptions are satisfied for $K=\mathbb{Q}(\sqrt{-1})$, under the Generalized Riemann Hypothesis (GRH).
Our strategy to prove Theorem A is to generalize Sharifi’s technique developed in [Sha02]. That is, we construct $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying the assumption of Theorem A in a group-theoretic way such that it generates $\mathrm{Gal}(\Omega/K(p^{\infty}))$, and we then use Conjecture 2.8 to conclude that the surjection $\mathrm{Gal}(\Omega/K(p^{\infty}))\to\mathrm{Gal}(\Omega^{\ast}/K(p^{\infty}))$ is an isomorphism.
To construct such elements, we need to introduce certain two-variable filtrations on the pro-$p$ geometric fundamental group $\Pi_{1,1}$ of $X$, on subgroups of the pro-$p$ mapping class group $\Gamma_{1,1}$ of type $(1,1)$ and on the absolute Galois group $G_{K}$, and establish their fundamental properties in Section 3.
As was mentioned in Remark 1.3 (1), the Galois group $\mathrm{Gal}(\Omega/K(p))$ has a nontrivial relation for every prime $p\geq 5$ which splits in $K$. This is one of the big differences from the previous result on the thrice-punctured projective line (cf. Remark 1.2), making it difficult to extend Sharifi’s approach to the case of once-punctured elliptic curves. Fortunately, however, the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a free pro-$p$ group (under the assumption of Theorem A), cf. Remark 4.3 (2). Therefore, once we obtain an explicit set of generators of $\mathrm{Gal}(\Omega/K(p^{\infty}))$, the rest of the proof follows in the same way as the previous result.
This paper proceeds as follows: In Section 2, we explain previous studies on the pro-$p$ outer Galois representations associated to once-punctured elliptic curves, especially the construction of a certain power series and Nakamura’s explicit formula for the power series. We then explain the properties of certain Kummer characters associated to the power series, propose an analogue of the Deligne-Ihara conjecture and state the main result in a precise manner. In Section 3, we introduce a certain two-variable version of the descending central series on various profinite groups and establish their fundamental properties. Section 4 is devoted to the proof of the main result.
Acknowledgement. This paper is a revised version of a part of the author’s doctoral thesis submitted to Kyoto University. The author sincerely thanks his advisor Akio Tamagawa for helpful comments, and Benjamin Collas for carefully reading drafts of the thesis. He also sincerely appreciates the referees for their constructive comments, which have greatly contributed to improving the content of this paper.
Notation
Indexes. For a pair $\boldsymbol{m}=(m_{1},m_{2})$ in $\mathbb{Z}^{2}$, we write $\lvert\boldsymbol{m}\rvert\coloneqq m_{1}+m_{2}$. Moreover, for an integer $w$, we write $\boldsymbol{m}\equiv 0\bmod w$ if $\boldsymbol{m}\in(w\mathbb{Z})^{2}$. For any two pairs $\boldsymbol{m}=(m_{1},m_{2})$ and $\boldsymbol{n}=(n_{1},n_{2})$, we write $\boldsymbol{m}\geq\boldsymbol{n}$ if $m_{i}\geq n_{i}$ holds for every $i=1,2$. Moreover, we write $\boldsymbol{m}>\boldsymbol{n}$ if $\boldsymbol{m}\geq\boldsymbol{n}$ and $\lvert\boldsymbol{m}\rvert>\lvert\boldsymbol{n}\rvert$. Finally, we often denote $(1,1)$ by $\boldsymbol{1}$.
Profinite Groups. For a profinite group $G$ and a subset $S$ of $G$, let $\langle S\rangle$ denote the closed subgroup of $G$ topologically generated by $S$. We say that $S$ generates $G$ if $G=\langle S\rangle$. Moreover, we say that $S$ strongly111 Note that the term “strong” is specific to this paper. In the book of Ribes-Zalesskii [RZ10, 2.4], the same notion is described as “$S$ is a generating set of $G$ converging to $1$”. generate $G$ if $S$ generates $G$ and $S$ converges to $1$, i.e. every open subgroup of $G$ contains all but a finite number of elements of $S$.
Let $\langle S\rangle_{{\rm normal}}$ denote the minimal normal closed subgroup of $G$ containing $S$. We call $\langle S\rangle_{{\rm normal}}$ the normal closure of $S$. We say that $S$ normally generates $G$ if $\langle S\rangle_{{\rm normal}}$ coincides with $G$.
The descending central series $\{G(m)\}_{m\geq 1}$ of $G$ is defined by
$G(1)\coloneqq G\quad\text{and}\quad G(m)\coloneqq\langle[G(m^{\prime}),G(m^{\prime\prime})]\mid m^{\prime}+m^{\prime\prime}=m\rangle\quad(m\geq 2),$
where, for every pair of subgroups $H,K\subset G$, the symbol $[H,K]$ denotes the topological closure of the commutator subgroup of $H$ and $K$. Note that the quotient $G^{\rm{ab}}=G/G(2)$ is the maximal abelian quotient of $G$, which we also denote by $G^{\rm{ab}}$. Moreover, for a prime number $p$, we denote the maximal pro-$p$ quotient of $G$ by $G^{(p)}$.
We denote the (continuous) automorphism group of $G$ by $\mathrm{Aut}(G)$, the inner automorphism group by $\mathrm{Inn}(G)$ and the outer automorphism group of $G$ by $\mathrm{Out}(G)\coloneqq\mathrm{Aut}(G)/\mathrm{Inn}(G)$. For $g\in G$, the associated inner automorphism is denoted by $\mathrm{inn}(g)$.
In this paper, we frequently use free pro-$p$ groups of countably infinite rank. We refer to Ribes-Zalesskii [RZ10, Lemma 3.3.4] for the characterization of free pro-$p$ groups on sets converging to $1$.
Number Fields. Throughout this paper, we work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ and a fixed embedding from $\bar{\mathbb{Q}}$ into $\mathbb{C}$. Every number field is considered to be a subfield of $\bar{\mathbb{Q}}$. For an integer $m\geq 1$, we denote the group of $m$-th roots of unity in $\bar{\mathbb{Q}}$ by $\mu_{m}$.
For a subfield $F$ of $\bar{\mathbb{Q}}$, we denote the absolute Galois group $\mathrm{Gal}(\bar{\mathbb{Q}}/F)$ of $F$ by $G_{F}$ and the ring of integers in $F$ by $O_{F}$. Moreover, for a place $v$ of $F$, we denote the $v$-adic completion of $F$ by $F_{v}$.
For an imaginary quadratic field $K$ and a nonzero integral ideal $\mathfrak{m}$ of $K$, we denote the ray class field of $K$ modulo $\mathfrak{m}$ by $K(\mathfrak{m})$ and write $K(\mathfrak{m}^{\infty})\coloneqq\cup_{n\geq 1}K(\mathfrak{m}^{n})$. If $\alpha\in O_{K}$ generates the ideal $\mathfrak{m}$, we sometimes denote $K(\mathfrak{m})$ and $K(\mathfrak{m}^{\infty})$ by $K(\alpha)$ and $K(\alpha^{\infty})$, respectively.
Elliptic Curves with Complex Multiplication (cf. Silverman [Sil94, Chapter II]). Let $K$ be an imaginary quadratic field of class number one, and $(E,O)$ an elliptic curve over $K$ with its origin $O\in E(K)$ having complex multiplication by $O_{K}$.
For an ideal $\mathfrak{m}$ of $O_{K}$, we denote the $\mathfrak{m}$-torsion subgroup scheme of $E$ by $E[\mathfrak{m}]$. The $G_{K}$-action on $E[\mathfrak{m}](\bar{\mathbb{Q}})$ determines an injective homomorphism
$\mathrm{Gal}(K(E[\mathfrak{m}])/K)\hookrightarrow\mathrm{Aut}(E[\mathfrak{m}](\bar{\mathbb{Q}}))\cong(O_{K}/\mathfrak{m})^{\times}.$
Moreover, this homomorphism induces an isomorphism
$\mathrm{Gal}(K(\mathfrak{m})/K)\xrightarrow{\sim}(O_{K}/\mathfrak{m})^{\times}/\mathrm{im}(O_{K}^{\times}),$
which does not depend on the choice of $E$.
[…]
$\chi_{1}\colon G_{K}\to\mathrm{Aut}(T_{\mathfrak{p}}(E))\cong\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon G_{K}\to\mathrm{Aut}(T_{\bar{\mathfrak{p}}}(E))\cong\mathbb{Z}_{p}^{\times}.$
Note that $\chi_{1}\chi_{2}=\chi_{\mathrm{cyc}}$, the $p$-adic cyclotomic character. For $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}^{2}$, we define
$\chi^{\boldsymbol{m}}\coloneqq\chi_{1}^{m_{1}}\chi_{2}^{m_{2}}\colon G_{K}\to\mathbb{Z}_{p}^{\times},$
which factors through $\mathrm{Gal}(K(p^{\infty})/K)$ if $m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert$. For a $\mathbb{Z}_{p}$-module $M$ on which the Galois group $G_{K}$ acts continuously, we denote the $\chi^{\boldsymbol{m}}$-twist of $M$ by $M(\boldsymbol{m})$. Using this notation, we have $T_{\mathfrak{p}}(E)\cong\mathbb{Z}_{p}(1,0)$, $T_{\bar{\mathfrak{p}}}(E)\cong\mathbb{Z}_{p}(0,1)$ and $\mathbb{Z}_{p}(m,m)$ is simply the $m$-th Tate twist $\mathbb{Z}_{p}(m)$.
2. Preliminaries
In this section, we prepare backgrounds which are necessary to state the main result of this paper (Theorem 2.12). In the following, we explain:
  • construction of a certain basis $\{x_{1},x_{2}\}$ of the pro-$p$ geometric fundamental group of once-punctured elliptic curve $X$, cf. Lemma 2.1,
  • construction of an elliptic analogue of Ihara’s universal power series for Jacobi sums [Iha86], and Nakamura’s explicit formula of that power series [Nak95], cf. Section 2.1,
  • conditional nonvanishing and surjectivity of certain Kummer characters associated to the power series [Ish25], cf. Theorem 2.5 and Theorem 2.6,
  • formulation of an analogue of the Deligne-Ihara conjecture, and state the main result, cf. Section 2.2.
In the following, let $K$ be an imaginary quadratic field of class number one, and $p\geq 5$ a prime which splits into two distinct primes in $O_{K}$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$. Let $(E,O)$ be an elliptic curve defined over $K$ having complex multiplication by $O_{K}$ with a Weierstrass form of $y^{2}=4x^{3}-g_{2}x-g_{3}$ with $g_{2},g_{3}\in K$. Let $L$ denote the period lattice corresponding to the fixed Weierstrass form.
We denote the once-punctured elliptic curve associated to $E$ by $X\coloneqq E-O$. We denote the pro-$p$ geometric fundamental group $\pi_{1}(\bar{X},\vec{O})^{(p)}$ with respect to the Weierstrass tangential basepoint at $O$ [Nak99, (2.7) Case 2] by $\Pi_{1,1}$, and identify $\Pi_{1,1}$ with a free pro-$p$ group of rank two with basis $\{\alpha,\beta\}$ in such a way that the commutator $[\alpha,\beta]$ generates an inertia subgroup at $O$.
Note that the pro-$p$ geometric fundamental group $\Pi_{1,0}$ of $E$ is isomorphic to $\Pi_{1,1}^{\rm{ab}}$ through a homomorphism $\Pi_{1,1}^{\rm{ab}}\to\Pi_{1,0}$ induced by the inclusion $X\subset E$. Moreover, there is a natural isomorphism $T_{p}(E)\xrightarrow{\sim}\Pi_{1,0}$. Throughout this paper, we work with the following basis $\{x_{1},x_{2}\}$ of $\Pi_{1,1}$, which is useful when considering the outer Galois action on $\Pi_{1,1}$:
Lemma 2.1.
There exists a basis $\{x_{1},x_{2}\}$ of $\Pi_{1,1}$ satisfying the following conditions:
  1. (1)
    Let $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) denote the image of $x_{1}$ (resp. $x_{2}$) in $\Pi_{1,1}^{\mathrm{ab}}\cong T_{p}(E)=T_{\mathfrak{p}}(E)\oplus T_{\bar{\mathfrak{p}}}(E)$. Then $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) is a basis of $T_{\mathfrak{p}}(E)$ (resp. $T_{\bar{\mathfrak{p}}}(E)$).
  2. (2)
    $z\coloneqq[x_{2},x_{1}]$ generates the inertia subgroup $\langle[\beta,\alpha]\rangle$.
Proof.
By a result of Kaneko [Kan89, Proposition 2], there exists an automorphism $\gamma$ of $\Pi_{1,1}$ such that the image of $\gamma(\alpha)$ (resp. $\gamma(\beta)$) in $\Pi_{1,1}^{\mathrm{ab}}$ generates $T_{\mathfrak{p}}(E)$ (resp. $T_{\bar{\mathfrak{p}}}(E)$) and the subgroup generated by $\gamma([\beta,\alpha])$ is conjugate to $\langle[\beta,\alpha]\rangle$. Since the inner automorphism group acts trivially on the maximal abelian quotient $\Pi_{1,1}^{\mathrm{ab}}$, we may assume that $\langle\gamma([\beta,\alpha])\rangle=\langle[\beta,\alpha]\rangle$. Now $x_{1}\coloneqq\gamma(\alpha)$ and $x_{2}\coloneqq\gamma(\beta)$ satisfy both conditions. ∎
2.1. Analogue of Ihara’s power series and Soulé characters
In this subsection, we introduce a certain power series which expresses the Galois action on the maximal metabelian quotient of $\Pi_{1,1}$, and Nakamura’s explicit formula of that power series. Then we introduce some non-vanishing results of the coefficients of the power series.
This power series is an elliptic analogue of the universal power series for Jacobi sums introduced by Ihara [Iha86], expressing the Galois action on the metabelian pro-$p$ fundamental group of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. We will not go into detail about the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$ here, and interested readers are encouraged to refer to the pioneering paper of Ihara [Iha86].
First, we define
$\tilde{\Gamma}_{1,1}\coloneqq\{f\in\mathrm{Aut}(\Pi_{1,1})\mid\text{$f$ preserves the conjugacy class of the inertia subgroup $\langle z\rangle$}\},$
its subgroup
$\Gamma_{1,1}^{{\dagger}}\coloneqq\{f\in\mathrm{Aut}(\Pi_{1,1})\mid\text{$f$ preserves $\langle z\rangle$}\},$
and a quotient $\Gamma_{1,1}\coloneqq\tilde{\Gamma}_{1,1}/\mathrm{Inn}(\Pi_{1,1})$ which is called the pro-$p$ mapping class group of type $(1,1)$. The group $\tilde{\Gamma}_{1,1}$ comes equipped with a descending central filtration $\{F^{m}\tilde{\Gamma}_{1,1}\}_{m\geq 1}$ called the weight filtration, which is defined by
$F^{m}\tilde{\Gamma}_{1,1}\coloneqq\mathrm{ker}\left(\tilde{\Gamma}_{1,1}\to\mathrm{Aut}(\Pi_{1,1}/\Pi_{1,1}(m+1))\right)$
for every $m\geq 1$. The weight filtration naturally induces filtrations on $\Gamma^{{\dagger}}_{1,1}$ and $\Gamma_{1,1}$ by
$F^{m}\Gamma^{{\dagger}}_{1,1}\coloneqq\Gamma^{{\dagger}}_{1,1}\cap F^{m}\tilde{\Gamma}_{1,1}\quad\text{and}\quad F^{m}\Gamma_{1,1}\coloneqq F^{m}\tilde{\Gamma}_{1,1}\mathrm{Inn}(\Pi_{1,1})/\mathrm{Inn}(\Pi_{1,1}),$
respectively.
Since the intersection $\cap_{m\geq 1}\Pi_{1,1}(m+1)$ of the descending central series is trivial, the intersection $\cap_{m\geq 1}F^{m}\tilde{\Gamma}_{1,1}$ is also trivial. Moreover, the intersection $\cap_{m\geq 1}F^{m}\Gamma_{1,1}$ is also trivial by a result of Asada [Asa95, Theorem 2].
Note that the normalizer subgroup of $\langle z\rangle$ in $\Pi_{1,1}$ is equal to $\langle z\rangle$ itself. It follows that the intersection $F^{m}\Gamma_{1,1}^{{\dagger}}\cap\mathrm{Inn}(\Pi_{1,1})$ is trivial for every $m\geq 3$ and is equal to $\langle\mathrm{inn}(z)\rangle$ for $m=1,2$. Therefore, the natural projection $F^{m}\Gamma_{1,1}^{{\dagger}}\to F^{m}\Gamma_{1,1}$ is an isomorphism for every $m\geq 3$ and, it is surjective with $\langle\mathrm{inn}(z)\rangle$ as its kernel for $m=1,2$ [Nak95, (4.4)].
We define two subgroups $\Psi^{{\dagger}}_{1}\subset\Psi^{{\dagger}}$ of $\mathrm{Aut}\left(\Pi_{1,1}/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\right)$ by
$\Psi^{{\dagger}}\coloneqq\{f\in\mathrm{Aut}\left(\Pi_{1,1}/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\right)\mid\text{$f$ preserves $\langle\bar{z}\rangle$}\}$
and
$\Psi^{{\dagger}}_{1}\coloneqq\ker\left(\Psi^{{\dagger}}\to\mathrm{Aut}(\Pi^{\rm{ab}})\right),$
where $\bar{z}\in\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ is the image of $z$ under the natural projection. We shall identify $\mathbb{Z}_{p}[[\Pi_{1,1}^{\rm{ab}}]]$ with $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ via $T_{i}\coloneqq x_{i}-1$ for $i=1,2$. The action of $\mathrm{Aut}(\Pi_{1,1}^{{\rm ab}})\cong\mathrm{GL}_{2}(\mathbb{Z}_{p})$ on $\Pi_{1,1}^{{\rm ab}}$ naturally extends to that of $\mathbb{Z}_{p}[[\Pi_{1,1}^{{\rm ab}}]]\cong\mathbb{Z}_{p}[[T_{1},T_{2}]]$.
Every element $f$ of $\Psi^{{\dagger}}_{1}$ is determined by a pair
$f(x_{i})x_{i}^{-1}\in\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\quad\text{for $i=1,2$}.$
Since $\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ is a free $\mathbb{Z}_{p}[[T_{1},T_{2}]]$-module generated by $\bar{z}$ [Iha86, Theorem 2], there exists a unique power series $G_{i}(f)\in\mathbb{Z}_{p}[[T_{1},T_{2}]]$ satisfying
$f(x_{i})x_{i}^{-1}=G_{i}(f)z\quad\text{for $i=1,2$}.$
Lemma 2.2 (Nakamura [Nak95, (4.7)], Tsunogai [Tsu95, Proposition 1.9]).
For every $f\in\Psi^{{\dagger}}_{1}$, two power series $G_{1}(f)$ and $G_{2}(f)$ satisfy
$T_{1}G_{2}(f)-G_{1}(f)T_{2}=0.$
If we write $H(f)\coloneqq\frac{G_{2}(f)}{T_{2}}=\frac{G_{1}(f)}{T_{1}}$, then $H\colon\Psi_{1}^{{\dagger}}\to\mathbb{Z}_{p}[[T_{1},T_{2}]]$ is an isomorphism. Moreover, the $\mathrm{GL}_{2}(\mathbb{Z}_{p})$-action on $\Psi_{1}^{{\dagger}}$ induced by the following exact sequence
$1\to\Psi_{1}^{{\dagger}}\to\Psi\to\mathrm{GL}_{2}(\mathbb{Z}_{p})\to 1$
makes $H\colon\Psi_{1}^{{\dagger}}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$ equivariant. Here, $(1)$ denotes the twist by the determinant character.
In the following, we shall exploit the section
$s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$
constructed by Nakamura [Nak95, (4.4)] as follows: First, by [Nak95, (4.4)], we have
$F^{1}\Gamma_{1,1}=F^{2}\Gamma_{1,1}=F^{3}\Gamma_{1,1}.$
Hence the image of $G_{K(E[p^{\infty}])}$ under $\rho_{X,p}$ is contained in the third component $F^{3}\Gamma_{1,1}$. By composing the inverse of the natural projection $F^{3}\Gamma_{1,1}^{{\dagger}}\xrightarrow{\sim}F^{3}\Gamma_{1,1}$, we obtain a homomorphism $s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$ which lifts $\rho_{X,p}\mid_{G_{K(E[p^{\infty}])}}$.
Definition 2.1.
We define a $\mathrm{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\alpha_{1,1}\colon G_{K(E[p^{\infty}])}\to\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$
as a composite of
  1. (1)
    the section $s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma_{1,1}^{{\dagger}}$ constructed above,
  2. (2)
    the natural projection $F^{3}\Gamma_{1,1}^{{\dagger}}\to\Psi^{{\dagger}}_{1}$, and
  3. (3)
    the isomorphism $H\colon\Psi^{{\dagger}}_{1}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$ constructed in Lemma 2.2.
In [Nak95], Nakamura obtained the following explicit formula of the power series $\alpha_{1,1}$ for general once-punctured elliptic curves. Note that we use the convention $0^{0}\coloneqq 1$ and regard $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ as a subring of $\mathbb{Q}_{p}[[U_{1},U_{2}]]$ where $U_{i}\coloneqq\log(1+T_{i})$ in the following formula.
Theorem 2.3 (Nakamura [Nak95, Theorem (A) and (3.11.5)]).
$\alpha_{1,1}(\sigma)=\sum_{m\geq 2\colon\text{even}}^{\infty}\frac{1}{1-p^{m}}\sum_{\begin{subarray}{c}\boldsymbol{m}=(m_{1},m_{2})\geq(0,0)\\ |\boldsymbol{m}|=m\end{subarray}}\kappa_{\boldsymbol{m}+\boldsymbol{1}}(\sigma)\frac{U_{1}^{m_{1}}U_{2}^{m_{2}}}{m_{1}!m_{2}!}$
holds for every $\sigma\in G_{K(E[p^{\infty}])}$, where $\kappa_{\boldsymbol{m}}\colon G_{K(E[p^{\infty}])}^{\rm{ab}}\to\mathbb{Z}_{p}$ is a Kummer character whose reduction modulo $p^{n}$ corresponds to the $p^{n}$-th roots of
$\prod_{\begin{subarray}{c}0\leq a,b<p^{n}\\ p\nmid\gcd(a,b)\end{subarray}}\theta(a\omega_{1,n}+b\omega_{2,n},L)^{a^{m_{1}-1}b^{m_{2}-1}}$
for every $n\geq 1$. Here $\theta(z,L)$ is the fundamental theta function defined in [Nak95, (2.1)], and the elements $(\omega_{1,n})_{n}\in T_{\mathfrak{p}}(E)$ and $(\omega_{2,n})_{n}\in T_{\bar{\mathfrak{p}}}(E)$ are defined as in Lemma 2.1 (1).
Remark 2.4.
Note that $\theta(z,L)$ is defined on $\mathbb{C}$, not on $E(\mathbb{C})$. Hence we have to choose a lift of $a\omega_{1,n}+b\omega_{2,n}\in E(\mathbb{C})\cong\mathbb{C}/L$ to $\mathbb{C}$, to consider the value $\theta(a\omega_{1,n}+b\omega_{2,n},L)$. However, by [Nak95, (2.2) Proposition (1)], the values of $\theta(z,L)$ at any such two lifts are equal up to $p$-th power roots of unity. Since $K(E[p^{\infty}])$ contains all $p$-th power roots of unity, the Kummer characters associated to $p$-th power roots of $\theta(z,L)$ at any such two lifts coincide with each other. This explains the notation in Theorem 2.3 above.
In the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$, it is known that coefficients of Ihara’s power series correspond to the Soulé characters $\{\chi_{m}\}_{m\geq 3,\mathrm{odd}}$ (see Ihara-Kaneko-Yukinari [IKY87, Theorem $\mathrm{A}_{2}$], for example). We therefore define their elliptic analogues as follows:
Definition 2.2 (The elliptic Soulé character).
For every $\boldsymbol{m}=(m_{1},m_{2})>\boldsymbol{1}$ such that $\lvert\boldsymbol{m}\rvert$ is even, we call the character $\kappa_{\boldsymbol{m}}$ appearing in Theorem 2.3 the $\boldsymbol{m}$-th elliptic Soulé character associated to $X$.
Nakamura proved that some linear combinations of the elliptic Soulé characters are nontrivial [Nak95, (3.12)]. In [Ish25], we conditionally proved that the elliptic Soulé characters arising from once-punctured CM elliptic curves are nontrivial. We briefly summarize previous results.
Since $\alpha_{1,1}$ is compatible with the action of $\mathrm{Gal}(K(E[p^{\infty}])/K)$, the $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ defines a $\mathrm{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\kappa_{\boldsymbol{m}}\colon G_{K(E[p^{\infty}])}^{\mathrm{ab}}\to\mathbb{Z}_{p}(\boldsymbol{m})$
for every $\boldsymbol{m}=(m_{1},m_{2})>\boldsymbol{1}$ such that $\lvert\boldsymbol{m}\rvert$ is even. Let
$I\coloneqq\{\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 1}^{2}\setminus\{\boldsymbol{1}\}\mid m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert\},$
and recall that $\Omega$ is the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
Theorem 2.5 ([Ish25, Theorem 1.5 (1) and Lemma 4.1]).
The $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ is trivial unless $\boldsymbol{m}$ is contained in $I$. For $\boldsymbol{m}\in I$, the character $\kappa_{\boldsymbol{m}}$ factors through $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ and, moreover, is nontrivial if $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite.
Here $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ denotes the second étale cohomology group of the spectrum of the ring of $p$-integers of $K$ with coefficient in $\mathbb{Z}_{p}(\boldsymbol{m})$, whose finiteness is a special case of a conjecture of Jannsen [Jan89, Conjecture 1]. For example, such a cohomology group is known to vanish for every $\boldsymbol{m}\in(p-1)\mathbb{Z}_{\geq 1}^{2}$ by e.g. [Ish25, Lemma 4.7]. It is also finite, when the coefficient is the $m$-th Tate twist $\mathbb{Z}_{p}(m)$ for some $m\geq 2$ [Sou79, page 287, Corollaire].
The following result gives a sufficient condition that certain elliptic Soulé characters are surjective:
Theorem 2.6 ([Ish25, Theorem 1.5]).
  1. (1)
    The character $\kappa_{\boldsymbol{m}}$ is not surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\geq(2,2)$ and $\boldsymbol{m}\equiv\boldsymbol{1}\mod p-1$.
  2. (2)
    If the class number of $K(p)$ is not divisible by $p$ and there are only two primes of $K(p)$ above $p$, then $\kappa_{\boldsymbol{m}}$ is surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\not\equiv\boldsymbol{1}\bmod p-1$.
Although $\kappa_{\boldsymbol{m}}$ is not surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\geq(2,2)$ and $\boldsymbol{m}\equiv\boldsymbol{1}\mod p-1$, we will see that it is nontrivial under the assumption of Theorem A, cf. Corollary A.3 in the appendix.
2.2. Analogue of the Deligne-Ihara conjecture and its consequence
In this subsection, we introduce a certain filtration on the absolute Galois group $G_{K}$ and formulate a conjecture on the structure of the graded Lie algebra associated to this filtration (Conjecture 2.8). This conjecture is an analogue of the so-called Deligne-Ihara conjecture in the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. We then study some fundamental properties of the fixed field of $\ker(\rho_{X,p})$ of the pro-$p$ outer Galois representation associated to $X$, and state the main result.
The absolute Galois group $G_{K}$ inherits a descending central filtration $\{F^{m}G_{K}\}_{m\geq 1}$ by $F^{m}G_{K}\coloneqq\rho_{X,p}^{-1}(F^{m}\Gamma_{1,1})$. Note that we have $F^{1}G_{K}=G_{K(E[p^{\infty}])}$. We form graded quotients and their directed sum as
$\mathfrak{g}_{m}\coloneqq F^{m}G_{K}/F^{m+1}G_{K}\text{ for $m\geq 1$ and }\mathfrak{g}\coloneqq\bigoplus_{m\geq 1}\mathfrak{g}_{m}.$
It is known that
  • Each $\mathfrak{g}_{m}$ is embedded into $F^{m}\Gamma_{1,1}/F^{m+1}\Gamma_{1,1}$, which is a free $\mathbb{Z}_{p}$-module of finite rank [NT93, Corollary (1.16), $(\mathrm{ii})$]. In particular, the $m$-th graded quotient $\mathfrak{g}_{m}$ is also a free $\mathbb{Z}_{p}$-module of finite rank.
  • Each $\mathfrak{g}_{m}\otimes\mathbb{Q}_{p}$ is isomorphic to a finite direct sum of (a finite direct sum of) $\mathbb{Q}_{p}(\boldsymbol{m})$’s as $G_{K}$-modules, where $\boldsymbol{m}\in\mathbb{Z}_{\geq 1}^{2}$ runs over indexes satisfying $|\boldsymbol{m}|=m$. This follows immediately from the $\mathrm{GL}_{2}(\mathbb{Z}_{p})$-equivariance of the commutative diagram given in [NT93, Theorem (1.14)].
We have the following lemma, relating the filtration with elliptic Soulé characters.
Lemma 2.7.
For $\boldsymbol{m}\in I$, the following assertions hold.
  1. (1)
    If $\kappa_{\boldsymbol{m}}\colon F^{1}G_{K}\to\mathbb{Z}_{p}(\boldsymbol{m})$ is nontrivial, then so is the character $\kappa_{\boldsymbol{m}}\mid_{F^{\lvert\boldsymbol{m}\rvert}G_{K}}$ restricted to ${F^{\lvert\boldsymbol{m}\rvert}G_{K}}$.
  2. (2)
    The character $\kappa_{\boldsymbol{m}}\mid_{F^{\lvert\boldsymbol{m}\rvert}G_{K}}$ factors through the $\lvert\boldsymbol{m}\rvert$-th graded quotient $\mathfrak{g}_{\lvert\boldsymbol{m}\rvert}$.
Proof.
In the following proof, we denote $\lvert\boldsymbol{m}\rvert$ by $m$, for simplicity. (1) Suppose that the restricted character $\kappa_{\boldsymbol{m}}\mid_{F^{m}G_{K}}$ is trivial. Then there exists an integer $1\leq n<m$ satisfying $\kappa_{\boldsymbol{m}}\mid_{F^{n+1}G_{K}}=0$ and $\kappa_{\boldsymbol{m}}\mid_{F^{n}G_{K}}\neq 0$. In particular, the $n$-th graded quotient $\mathfrak{g}_{n}\otimes\mathbb{Q}_{p}$ has a nontrivial $\chi^{\boldsymbol{m}}$-isotypic component, which is absurd since $n<m=\lvert\boldsymbol{m}\rvert$.
(2) It suffices to prove that the character $\kappa_{\boldsymbol{m}}$ vanishes when restricted to $F^{m+2}G_{K}$, since we have $F^{m+1}G_{K}=F^{m+2}G_{K}$ [Nak95, (4.2) Proposition]. By the construction of the power series $\alpha_{1,1}(\sigma)\in\mathbb{Z}_{p}[[T_{1},T_{2}]]$, it follows that, for every $\sigma\in F^{m+2}G_{K}$,
$T_{1}\alpha_{1,1}(\sigma)z\in\Pi_{1,1}(m+3)\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\subset\mathbb{Z}_{p}[[T_{1},T_{2}]]z.$
Note that the submodule $\Pi_{1,1}(m+3)\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ corresponds to the ideal $J^{m+1}$, where $J$ denotes the augmentation ideal of $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ [Iha86, (19) on page 67]. It follows that the power series $T_{1}\alpha_{1,1}(\sigma)$ is contained in $J^{m+1}$, and hence $\alpha_{1,1}(\sigma)\in J^{m}$. Hence every coefficient of $\alpha_{1,1}(\sigma)$ of a monomial of total degree less than $m$ vanishes. By observing Theorem 2.3, it follows that $\kappa_{\boldsymbol{n}}(\sigma)$ vanishes for every $\boldsymbol{n}\in I$ such that $\lvert\boldsymbol{n}\rvert\leq m-2$, as desired. ∎
The Deligne-Ihara conjecture [Sha02] states that a $\mathbb{Q}_{p}$-graded Lie algebra, defined in a similar manner to $\mathfrak{g}\otimes\mathbb{Q}_{p}$ using the pro-$p$ outer Galois representation of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$, is a free graded Lie algebra in each odd degree $m\geq 3$. We formulate an analogue of this conjecture, as follows:
Conjecture 2.8.
For every $\boldsymbol{m}\in I$, choose $\sigma_{\boldsymbol{m}}$ such that
  1. (1)
    The element $\sigma_{\boldsymbol{m}}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component of $\mathfrak{g}_{\boldsymbol{\lvert m\rvert}}$, and
  2. (2)
    The image $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})\in\mathbb{Z}_{p}(\boldsymbol{m})$ under $\kappa_{\boldsymbol{m}}$ is nonzero, and generates $\kappa_{\boldsymbol{m}}(\mathfrak{g}_{|\boldsymbol{m}|}).$
Then the graded Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$ is freely generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$.
Remark 2.9.
(1) Using the theory of weighted completions of profinite groups developed by Hain-Matsumoto [HM03], one can prove the generation part of Conjecture 2.8, assuming that $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite for every $\boldsymbol{m}\in I$. At the writing of this paper, however, we do not know how to prove the freeness portion of the conjecture. We hope to discuss the structure of $\mathfrak{g}\otimes\mathbb{Q}_{p}$ in future research.
(2) For the second assumption of the Conjecture 2.8, it is enough to only assume $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})\neq 0$, since we concentrated on the structure of the $\mathbb{Q}_{p}$-graded Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$. However, the latter condition becomes essential when considering the structure of the $\mathbb{Z}_{p}$-graded Lie algebra $\mathfrak{g}$. Our proof of Theorem A shows that, under the assumption of Theorem A, one can choose the set $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying two assumptions of Conjecture 2.8 so that it even freely generates $\mathfrak{g}$.
We turn our attention to field-theoretic properties of the fixed field of the kernel $\ker(\rho_{X,p})$ of the pro-$p$ outer Galois representation of the once-punctured elliptic curve $X=E-O$.
Lemma 2.10.
The fixed field $\bar{K}^{\ker(\rho_{X,p})}$ is a pro-$p$ extension of $K(E[p])$ unramified outside $p$.
Proof.
It suffices to show that $X\times_{K}K(E[p])$ has good reduction everywhere outside $p$, by virtue of specialization homomorphisms of étale fundamental groups and the fact that $\ker\left(\mathrm{Out}(\Pi_{1,1})\to\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)\right)$ is pro-$p$. By the Néron-Ogg-Shafarevich criterion, it suffices to show that every inertia subgroup $I$ associated to a prime of $K(E[p])$ outside $p$ acts trivially on $T_{p}(E)$.
Since $E$ has complex multiplication, it has everywhere potentially good reduction. Hence the image of $I$ in $\mathrm{Aut}(T_{p}(E))$ is finite. Moreover, the image of $I$ is contained in the subgroup $\ker(\mathrm{GL}(T_{p}(E))\to\mathrm{GL}(E[p]))$, which is a torsion-free pro-$p$ group for every $p\geq 5$. Hence the image of $I$ is trivial. ∎
Lemma 2.11.
The field $\bar{K}^{\ker(\rho_{X,p})}$ can be written as a compositum of $K(E[p])$ and a subfield $\Omega^{\ast}\subset\Omega$, in such a way that $\mathrm{Gal}(\bar{K}^{\ker(\rho_{X,p})}/K(p))$ splits into the direct product of $\mathrm{Gal}(K(E[p])/K(p))$ and $\mathrm{Gal}(\Omega^{\ast}/K(p))$.
Proof.
Consider the exact sequence
$1\to\rho_{X,p}(G_{K(E[p])})\to\rho_{X,p}(G_{K(p)})\to\mathrm{Gal}(K(E[p])/K(p))\to 1$
Since $\rho_{X,p}(G_{K(E[p])})$ is a pro-$p$ group and $\mathrm{Gal}(K(E[p])/K(p))$ is a finite prime-to-$p$ group (here we use $p\geq 5$), this sequence splits by the Schur-Zassenhaus theorem [Ser97, 5.9, Corollary 1]. Let $t\colon\mathrm{Gal}(K(E[p])/K(p))\to\rho_{X,p}(G_{K(p)})$ be an arbitrary section.
There is a natural homomorphism $\mathrm{Aut}_{K}(X)\to\mathrm{Out}(\Pi_{1,1})$x, and the image of $\mathrm{Aut}_{K}(X)$ centralizes $\rho_{X,p}(G_{K})$.
Through the $G_{K}$-action on $E[p]$, the Galois group $\mathrm{Gal}(K(E[p])/K(p))$ is identified with a subgroup of the image of $\mathrm{Aut}_{K}(X)$ in $\mathrm{Aut}_{O_{K}}(E[p])=(O_{K}/p)^{\times}$. Moreover $\mathrm{Aut}_{K}(X)=O_{K}^{\times}$ injects into $(O_{K}/p)^{\times}$ since $p$ is prime to the order of $O_{K}^{\times}$. Hence for every $g\in\mathrm{Gal}(K(E[p])/K(p))$, there exists a unique $\tilde{g}\in\mathrm{Aut}_{K}(X)$ such that $t(g)$ and $\tilde{g}$ coincide in $\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)=\mathrm{Aut}(E[p])$. Since $t(g)\tilde{g}^{-1}$ has a prime-to-$p$ order and is contained in $\ker\left(\mathrm{Out}(\Pi_{1,1})\to\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)\right)$ which is pro-$p$, we conclude $t(g)=\tilde{g}$.
It shows that the image of the section $t$ coincides with a subgroup of $\mathrm{Aut}_{K}(X)$, which is contained in the center of $\rho_{X,p}(G_{K})$. In particular, the section $t$ induces the decomposition
$\rho_{X,p}(G_{K(p)})=\rho_{X,p}(G_{K(E[p])})\times\mathrm{imz}(t).$
Let $\Omega^{\ast}$ be the field corresponding to the kernel of the projection $\rho_{X,p}(G_{K(p)})\to\rho_{X,p}(G_{K(E[p])})$. It is clear that $\Omega^{\ast}$ is a pro-$p$ extension of $K(p)$ unramified outside $p$, and
$\rho_{X,p}(G_{K(p)})=\mathrm{Gal}(\bar{K}^{\ker(\rho_{X,p})}/K(p))$
admits the required decomposition. ∎
Lemma 2.11 allows us to formulate the following question, which is an analogue of Anderson-Ihara’s question in Introduction:
Is the field $\Omega^{\ast}$ is equal to $\Omega$?
Recalling Sharifi’s result, one may wonder if this question has also an affirmative answer, assuming that Conjecture 2.8 holds and that a certain arithmetic condition on $p$ is satisfied. This is exactly what we prove in the following:
Theorem 2.12.
Let $p\geq 5$ be a prime which splits in $K$, and assume that the following conditions hold:
  1. (1)
    the class number of $K(p)$ is not divisible by $p$,
  2. (2)
    there are exactly two primes of $K(p^{\infty})$ above $p$, and
  3. (3)
    Conjecture 2.8 holds.
Then the field $\Omega^{\ast}$ coincides with $\Omega$.
3. Two-variable filtrations on profinite groups
In this section, we define two-variable filtrations on various profinite groups, e.g. the pro-$p$ fundamental groups of once-punctured elliptic curves, subgroups of the pro-$p$ mapping class group of type $(1,1)$ and Galois groups, and establish some fundamental properties. Throughout this section, let $\Pi$ denote a free pro-$p$ group of rank two with basis $\{x,y\}$ and define $z\coloneqq[y,x]$.
3.1. Two-variable filtration on free pro-$p$ group of rank two
First, we define a two-variable filtration on $\Pi$, which is an analogue of the descending central series.
Definition 3.1.
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we inductively define a normal subgroup $\Pi(\boldsymbol{m})$ of $\Pi$ as follows:
  1. (1)
    Let $\Pi(1,0)$ (resp. $\Pi(0,1)$) be the normal closure of $x$ (resp. $y$) in $\Pi$.
  2. (2)
    For $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}$ with $|\boldsymbol{m}|\geq 2$, we define the subgroup $\Pi(\boldsymbol{m})\subset\Pi$ by
    $\Pi(\boldsymbol{m})\coloneqq\langle[\Pi(\boldsymbol{m}^{\prime}),\Pi(\boldsymbol{m}^{\prime\prime})]\mid\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}\text{ where $\boldsymbol{m}^{\prime},\boldsymbol{m}^{\prime\prime}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$}\rangle.$
    Since $\Pi(1,0)$ and $\Pi(0,1)$ are normal subgroups, the subgroup $\Pi(\boldsymbol{m})$ is also normal.
The definition of the two-variable filtration depends on the choice of the basis $\{x,y\}$ of $\Pi$. More precisely, $\Pi(1,0)$ (resp. $\Pi(0,1)$) depends on the image $\bar{y}$ of $y$ (resp. $\bar{x}$ of $x$) in the maximal abelian quotient $\Pi^{\mathrm{ab}}$ of $\Pi$.
Example 3.1.
By definition, $\Pi(1,1)$ coincides with the commutator subgroup $\Pi(2)$.
Remark 3.1.
Let $\mathcal{L}$ be a free graded Lie algebra of rank two generated by $\{x_{\mathcal{L}},y_{\mathcal{L}}\}$. The filtration given in Definition 3.1 may be regarded as a group-theoretic analogue of the bi-grading on $\mathcal{L}$ induced by degrees with respect to $x_{\mathcal{L}}$ and $y_{\mathcal{L}}$.
We have the following inclusions and equalities:
Lemma 3.2.
Let $\boldsymbol{m}=(m_{1},m_{2}),\boldsymbol{n}=(n_{1},n_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$ and $m\geq 2$.
(1) The inclusion $\Pi(\boldsymbol{m})\subset\Pi(|\boldsymbol{m}|)$ holds.
(2) The equality $\Pi(m,0)=\Pi(m,1)$ holds. Similarly, we have $\Pi(0,m)=\Pi(1,m)$.
(3) If $\boldsymbol{m}\geq\boldsymbol{n}$, then the inclusion $\Pi(\boldsymbol{m})\subset\Pi(\boldsymbol{n})$ holds.
Proof.
(1) The assertion immediately follows by induction on $\lvert\boldsymbol{m}\rvert$.
(2) The inclusion $\Pi(m,1)\subset\Pi(m,0)$ follows by induction on $m$. To prove the opposite inclusion by induction on $m$, it suffices to prove the assertion for $m=2$. Note that the group $\Pi(2,0)/\Pi(2,1)$ is normally generated by the image of the commutator map
$\Pi(1,0)/\Pi(\boldsymbol{1})\times\Pi(1,0)/\Pi(\boldsymbol{1})\to\Pi(2,0)/\Pi(2,1).$
However, since $\Pi(1,0)/\Pi(\boldsymbol{1})$ is generated by the image of $x$, the image of the commutator map is trivial. Hence the equality $\Pi(2,0)=\Pi(2,1)$ follows.
(3) We prove the assertion by induction on $|\boldsymbol{m}|+|\boldsymbol{n}|$. If $|\boldsymbol{m}|+|\boldsymbol{n}|=2$, then the assertion is clear. Let us assume that $|\boldsymbol{m}|+|\boldsymbol{n}|>2$. Since $\Pi(\boldsymbol{m})$ is normally generated by $[\Pi(\boldsymbol{m}^{\prime}),\Pi(\boldsymbol{m}^{\prime\prime})]$ with $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$, the assertion follows if there exist $\boldsymbol{n}^{\prime}$ and $\boldsymbol{n}^{\prime\prime}$ such that $\boldsymbol{m}^{\prime}\geq\boldsymbol{n}^{\prime}$, $\boldsymbol{m}^{\prime\prime}\geq\boldsymbol{n}^{\prime\prime}$ and $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$, by induction hypothesis. Such a pair $(\boldsymbol{n}^{\prime},\boldsymbol{n}^{\prime\prime})$ clearly exists unless $(m_{1},n_{1})$ or $(m_{2},n_{2})$ is equal to $(1,0)$. However, the assertion in this exceptional case also follows by using (2). ∎
Definition 3.2.
For $\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define graded quotients $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$ of $\Pi$ as
$\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(1,0))\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(0,1)).$
Note that $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$) is a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$ (resp. $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$)-module where the group $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$) acts by conjugation.
Example 3.2.
(1) $\mathrm{gr}^{(1,0)}_{1}\Pi=\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module of rank one generated by $x$, cf. [Iha86, Theorem 2.2]. Similarly, $\mathrm{gr}^{(0,1)}_{2}\Pi$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$-module of rank one generated by $y$.
(2) $\mathrm{gr}^{(1,0)}_{2}\Pi=\Pi(1,0)/\Pi(\boldsymbol{1})$ is a free $\mathbb{Z}_{p}$-module of rank one generated by the image of $x$ on which $\Pi/\Pi(1,0)$ acts trivially. Similarly, $\mathrm{gr}^{(0,1)}_{1}\Pi$ is generated by $y$ on which $\Pi/\Pi(0,1)$ acts trivially. On the other hand, by Lemma 3.2(2), we have $\mathrm{gr}^{(m,0)}_{2}\Pi=\mathrm{gr}^{(0,m)}_{1}\Pi=0$ for every $m\geq 2$.
3.2. Two-variable filtration on pro-$p$ mapping class group
In this subsection, we define a two-variable filtration on subgroups of the automorphism group of $\Pi$. The aim of this subsection is to prove Proposition 3.9, which computes the action on the graded quotients associated to the two-variable filtration. We note that the content of this subsection may be regarded as a two-variable variant of a part of a work of Nakamura-Tsunogai [NT93], studying the structure of graded quotients of pro-$p$ mapping class groups with respect to weight filtrations.
Definition 3.3.
We define two subgroups $\tilde{\Gamma}$ and $\Gamma^{{\dagger}}$ of $\mathrm{Aut}(\Pi)$ as
$\tilde{\Gamma}\coloneqq\left\{f\in\mathrm{Aut}(\Pi)\,\middle|\,\begin{aligned}&\text{$\bar{f}$ preserves $\langle\bar{x}\rangle$ and $\langle\bar{y}\rangle$ respectively, and}\\ &\text{$f$ preserves the conjugacy class of $\langle z\rangle$}\end{aligned}\right\}\text{and}$
$\Gamma^{{\dagger}}\coloneqq\{f\in\tilde{\Gamma}\mid\text{$f$ preserves $\langle z\rangle$}\},$
where $\bar{f}\in\mathrm{Aut}(\Pi^{\mathrm{ab}})$ is the automorphism of $\Pi$ induced by $f$ in. Moreover, we set $\Gamma\coloneqq\tilde{\Gamma}/\mathrm{Inn}(\Pi)$, which is a subgroup of $\mathrm{Out}(\Pi)$.
One can easily observe that the subgroup $\Pi(\boldsymbol{m})$ defined in the last subsection is preserved under the action of $\tilde{\Gamma}$. By the definition of $\tilde{\Gamma}$, there exist two homomorphisms [^3] This notation is ambiguous since we already use the symbols $\chi_{1}$ and $\chi_{2}$ to indicate the characters $G_{K}\to\mathbb{Z}_{p}^{\times}$. However, if we identify $\Pi$ with the fundamental group $\Pi_{1,1}$ via $x=x_{1}$ and $y=x_{2}$, then the image of the pro-$p$ outer Galois representation $\rho_{X,p}$ is contained in $\Gamma$, and these notations become compatible.
$\chi_{1}\colon\tilde{\Gamma}\to\mathrm{Aut}(\langle\bar{x}\rangle)=\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon\tilde{\Gamma}\to\mathrm{Aut}(\langle\bar{y}\rangle)=\mathbb{Z}_{p}^{\times}$
which are characterized by the equalities $\bar{f}(\bar{x})=\bar{x}^{\chi_{1}(f)}$ and $\bar{f}(\bar{y})=\bar{y}^{\chi_{2}(f)}$ for every $f\in\tilde{\Gamma}$, respectively. Note that these characters factor through the quotient $\Gamma$. There is also a homomorphism $\Gamma^{{\dagger}}\to\mathrm{Aut}(\langle z\rangle)=\mathbb{Z}_{p}^{\times}$ by the definition of $\Gamma^{{\dagger}}$, which is found to coincide with the product $\chi_{1}\chi_{2}$ since the commutator map $\Pi/\Pi(2)\times\Pi/\Pi(2)\to\Pi(2)/\Pi(3)$ is bilinear.
Remark 3.3.
If we identify $\Pi$ with the pro-$p$ fundamental group $\Pi_{1,1}$ of the once-punctured CM elliptic curve $X=E-O$ via $x=x_{1}$ and $y=x_{2}$, then the group $\tilde{\Gamma}$ (resp. $\Gamma^{{\dagger}}$, $\Gamma$) is a subgroup of $\tilde{\Gamma}_{1,1}$ (resp. $\Gamma_{1,1}^{{\dagger}}$, $\Gamma_{1,1}$) introduced in Section 2.1. Moreover, through this identification, the action of the absolute Galois group $G_{K}$ preserves $\langle\bar{x}\rangle$ and $\langle\bar{y}\rangle$, since they are isomorphic to the $\mathfrak{p}$-adic and the $\bar{\mathfrak{p}}$-adic Tate module of $E$, respectively. In particular, the image of the pro-$p$ outer Galois representation $\rho_{X,p}:G_{K}\to\Gamma_{1,1}$ is contained in $\Gamma$.
Recall that, in Section 2.1, the weight filtration $\{F^{m}\tilde{\Gamma}\}_{m\geq 1}$ on $\tilde{\Gamma}$ is defined by
$F^{m}\tilde{\Gamma}\coloneqq\ker\left(\tilde{\Gamma}\to\mathrm{Aut}(\Pi/\Pi(m+1))\right)$
for $m\geq 1$. By replacing the descending central series with the two-variable filtration $\{\Pi(\boldsymbol{m})\}_{\boldsymbol{m}}$ on $\Pi$, we define a two-variable analogue of the weight filtration as follows.
Definition 3.4.
For every $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define subgroups $F^{\boldsymbol{m}}\tilde{\Gamma}$, $F^{\boldsymbol{m}}\Gamma^{{\dagger}}$ and $F^{\boldsymbol{m}}\Gamma$ by
$F^{\boldsymbol{m}}\tilde{\Gamma}\coloneqq\ker\left(\tilde{\Gamma}\to\prod_{\boldsymbol{k}\in\{(0,1),(1,0)\}}\mathrm{Aut}\left(\Pi(\boldsymbol{k})/\Pi(\boldsymbol{m}+\boldsymbol{k})\right)\right),$
$F^{\boldsymbol{m}}\Gamma^{{\dagger}}\coloneqq F^{\boldsymbol{m}}\Gamma\cap\Gamma^{{\dagger}},\quad\text{and}\quad F^{\boldsymbol{m}}\Gamma\coloneqq\mathrm{Im}(F^{\boldsymbol{m}}\tilde{\Gamma}\to\Gamma).$
Note that, since $[\Pi(\boldsymbol{m}),\Pi(\boldsymbol{k})]\subset\Pi(\boldsymbol{m}+\boldsymbol{k})$ for $\boldsymbol{k}\in\{(1,0),(0,1)\}$, the inner automorphism group $\mathrm{Inn}_{\Pi(\boldsymbol{m})}(\Pi)$ of $\Pi$ induced by elements of $\Pi(\boldsymbol{m})$ is contained in $F^{\boldsymbol{m}}\tilde{\Gamma}$.
Note that the subgroup $F^{\boldsymbol{m}}\tilde{\Gamma}$ acts trivially on the quotient $\Pi/\left(\Pi(\boldsymbol{m})\cap\Pi(|\boldsymbol{m}|+1)\right)$ by definition and Lemma 3.2, hence on $\Pi/\Pi(|\boldsymbol{m}|+1)$. We therefore obtain the following comparison with the usual weight filtration:
$F^{\boldsymbol{m}}\tilde{\Gamma}\subset F^{|\boldsymbol{m}|}\tilde{\Gamma}.$
The point of Definition 3.4 is to impose conditions on the images of $x$ and $y$ separately, i.e. $f(x)x^{-1}\in\Pi(\boldsymbol{m}+(1,0))$ and $f(y)y^{-1}\in\Pi(\boldsymbol{m}+(0,1))$ for each $f\in F^{\boldsymbol{m}}\tilde{\Gamma}$. In fact, the following lemma characterizes elements of $F^{\boldsymbol{m}}\tilde{\Gamma}$:
Lemma 3.4.
Let $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$ and $f\in\tilde{\Gamma}$. Then the automorphism $f$ is contained in $F^{\boldsymbol{m}}\tilde{\Gamma}$ if and only if $f(x)x^{-1}\in\Pi(\boldsymbol{m}+(1,0))$ and $f(y)y^{-1}\in\Pi(\boldsymbol{m}+(0,1))$.
Proof.
It suffices to prove only the ‘‘if’’ part of the assertion. Since the group $\Pi(1,0)$ is generated by $\{y^{n}xy^{-n}\}_{n\geq 0}$, it suffices to prove that $f(y^{n}xy^{-n})y^{n}x^{-1}y^{-n}\in\Pi(\boldsymbol{m}+(1,0))$ for every $n\geq 0$. We compute this term as follows.
$f(y^{n}xy^{-n})y^{n}x^{-1}y^{-n}$ $=f(y)^{n}f(x)f(y)^{-n}y^{n}x^{-1}y^{-n}$
$\equiv f(y)^{n}f(y)^{-n}y^{n}f(x)x^{-1}y^{-n}\bmod\Pi(\boldsymbol{m}+(1,0))$
$=y^{n}f(x)x^{-1}y^{-n}\equiv 1.$
Here, we use $[f(y)^{-n}y^{n},f(x)]\in[\Pi(\boldsymbol{m}+(0,1)),\Pi(1,0)]\subset\Pi(\boldsymbol{m}+\boldsymbol{1})\subset\Pi(\boldsymbol{m}+(1,0))$ to establish the first congruence. A similar computation shows that $f$ also acts trivially on $\Pi(0,1)/\Pi(\boldsymbol{m}+(0,1))$. ∎
Moreover, we have the following lemma:
Lemma 3.5.
Let $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then the group $F^{\boldsymbol{m}}\tilde{\Gamma}$ acts trivially on the quotient $\Pi(\boldsymbol{n})/\Pi(\boldsymbol{n}+\boldsymbol{m})$ for every $\boldsymbol{n}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$.
Proof.
We prove the assertion by induction on $\lvert\boldsymbol{n}\rvert$. First, if $\lvert\boldsymbol{n}\rvert=1$, the assertion follows by the definition of $F^{\boldsymbol{m}}\tilde{\Gamma}$. In general, since $\Pi(\boldsymbol{n})$ is generated by subgroups $[\Pi(\boldsymbol{n}^{\prime}),\Pi(\boldsymbol{n}^{\prime\prime})]$ such that $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$ and $0<|\boldsymbol{n}^{\prime}|,|\boldsymbol{n}^{\prime\prime}|<|\boldsymbol{n}|$, the assertion follows by the fact that the commutator map
$\Pi(\boldsymbol{n}^{\prime})/\Pi(\boldsymbol{n}^{\prime}+\boldsymbol{m})\times\Pi(\boldsymbol{n}^{\prime\prime})/\Pi(\boldsymbol{n}^{\prime}+\boldsymbol{m})\to\Pi(\boldsymbol{n})/\Pi(\boldsymbol{n}+\boldsymbol{m})$
is compatible with the action of $\tilde{\Gamma}$. ∎
Similar to the case of $\Pi$, we consider two-variable graded quotients as follows.
Definition 3.5.
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define $\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$ by
$\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(1,0)}\tilde{\Gamma}\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(0,1)}\tilde{\Gamma}.$
Similarly, we define $\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma$ by
$\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(1,0)}\Gamma\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(0,1)}\Gamma.$
Note that, by definition, we have natural surjective homomorphisms
$\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma.$
The following homomorphisms are useful in studying these quotients.
Definition 3.6.
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define homomorphisms $i_{\boldsymbol{m},1}$ and $i_{\boldsymbol{m},2}$ by
$i_{\boldsymbol{m},1}\colon\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1})\quad\text{and}$
$i_{\boldsymbol{m},2}\colon\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1}).$
Note that Lemma 3.4 implies that both $i_{\boldsymbol{m},1}$ and $i_{\boldsymbol{m},2}$ are injective.
Remark 3.6.
Two homomorphisms $i_{\boldsymbol{m},1}$ and $i_{\boldsymbol{m},2}$ defined above are analogues of the homomorphism $\tilde{\iota}_{m}$ defined in [NT93, (1.11.2)], which relates the (one-variable) graded quotients of the pro-$p$ mapping class groups with those of the pro-$p$ fundamental groups.
Write
$\tilde{\Gamma}_{1}\coloneqq\{\gamma\in\tilde{\Gamma}\mid\gamma(y)y^{-1}\in\Pi(2)\}\quad\text{and}\quad\tilde{\Gamma}_{2}\coloneqq\{\gamma\in\tilde{\Gamma}\mid\gamma(x)x^{-1}\in\Pi(2)\}.$
Moreover, we define
$\Gamma^{{\dagger}}_{1}\coloneqq\tilde{\Gamma}_{1}\cap\Gamma^{{\dagger}},\quad\Gamma_{1}\coloneqq\mathrm{Im}(\tilde{\Gamma}_{1}\to\Gamma),$
$\Gamma^{{\dagger}}_{2}\coloneqq\tilde{\Gamma}_{2}\cap\Gamma^{{\dagger}},\quad\Gamma_{2}\coloneqq\mathrm{Im}(\tilde{\Gamma}_{2}\to\Gamma).$
Then the action of $\tilde{\Gamma}_{1}/F^{(1,0)}\tilde{\Gamma}$ (resp. $\tilde{\Gamma}_{2}/F^{(0,1)}\tilde{\Gamma}$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$) commutes with the action of $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$).
In the following, we study the action of the quotient $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}$) of the groups $\Pi$, $\tilde{\Gamma}$ and $\Gamma$.
Lemma 3.7.
For every $\gamma\in\Gamma^{{\dagger}}_{1}$, we have $\gamma(x)x^{-\chi_{1}(\gamma)}\in\Pi(2,0)$. Similarly, the element $\gamma(y)y^{-\chi_{2}(\gamma)}$ is contained in $\Pi(0,2)$ for every $\gamma\in\Gamma^{{\dagger}}_{2}$.
Proof.
We only prove the first assertion, since the proof of the second one is similar. Since $\gamma\in\Gamma^{{\dagger}}_{1}$ acts $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linearly on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by $x$, the action of $\gamma$ is given by a scalar multiplication by an element of $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]^{\times}$. However, since we have $z=[y,x]=(y-1)x\neq 0$ in $\Pi(1,0)/\Pi(2,0)$ and $\gamma(z)=z^{\chi_{1}(\gamma)}$, such a scalar coincides with $\chi_{1}(\gamma)$. Hence we have $\gamma(x)\equiv x^{\chi_{1}(\gamma)}\bmod\Pi(2,0)$ as desired. ∎
Lemma 3.8.
Let $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then the action of the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ on the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$ on $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$ is given by $\chi_{2}^{m_{2}}$.
Proof.
Fix an element $\gamma\in\Gamma^{{\dagger}}_{1}$, and we prove by induction on $|\boldsymbol{m}|$ that the action of $\gamma$ on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is given by multiplication by $\chi_{1}(\gamma)^{m_{1}}$. If $|\boldsymbol{m}|=1$, the assertion follows from (the proof of) Lemma 3.7. In general, the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is generated (as a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module) by the image of commutator maps
$[\;\cdot\;,\;\cdot\;]\colon\mathrm{gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\mathrm{gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi\to\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi,$
where $\boldsymbol{m}^{\prime}=(m_{1}^{\prime},m_{2}^{\prime})$ and $\boldsymbol{m}^{\prime\prime}=(m_{1}^{\prime\prime},m_{2}^{\prime\prime})$ satisfy $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$. Since this pairing is bilinear, for every $(\tau^{\prime},\tau^{\prime\prime})\in\mathrm{gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\mathrm{gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi$, we have
$\gamma([\tau^{\prime},\tau^{\prime\prime}])=[\gamma(\tau^{\prime}),\gamma(\tau^{\prime\prime})]=[\chi_{1}^{m^{\prime}_{1}}(f)\tau^{\prime},\chi_{1}^{m^{\prime\prime}_{1}}(\gamma)\tau^{\prime\prime}]=\chi_{1}^{m_{1}}(\gamma)[\tau^{\prime},\tau^{\prime\prime}]$
by using the induction hypothesis. This concludes the proof. ∎
We now prove the main result in this subsection.
Proposition 3.9.
For every $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, the homomorphism
$i_{\boldsymbol{m},1}\colon\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi$
is compatible with the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$. Here,x $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})$ denotes the $\chi_{1}^{-1}$-twist of $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi$. Similarly, the homomorphism
$i_{\boldsymbol{m},2}\colon\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi(\chi_{2}^{-1})$
is compatible with the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$. In particular, the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
Proof.
Before proving the assertion, we note that the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on the graded quotients $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ by the character $\chi_{1}^{m_{1}}$, by Lemma 3.8. Take an arbitrary element $\gamma$ of $\Gamma^{{\dagger}}_{1}$. We prove the first assertion by computing the image of $\gamma\cdot f$ under the map $i_{\boldsymbol{m},1}$ for an arbitrary $f\in F^{\boldsymbol{m}}\tilde{\Gamma}$.
First, we have
$(\gamma\cdot f)(y)y^{-1}$ $=(\gamma f\gamma^{-1})(y)y^{-1}$
$=\gamma(f(\gamma^{-1}(y))\gamma^{-1}(y^{-1})).$
Since there exists $\alpha\coloneqq y^{-1}\gamma^{-1}(y)$ is contained in $\Pi(2)=\Pi(\boldsymbol{1})$, we have
$\gamma(f(\gamma^{-1}(y))\gamma^{-1}(y^{-1}))=\gamma(f(y\alpha)\alpha^{-1}y^{-1})\equiv\gamma(f(y)y^{-1})\bmod\Pi(\boldsymbol{m}+\boldsymbol{1}).$
Here, we use $f(\alpha)\alpha^{-1}\in\Pi(\boldsymbol{m}+\boldsymbol{1})$, which follows from Lemma 3.5, to deduce the last congruence. By Lemma 3.8, the last term is equal to $\chi_{1}^{m_{1}}(\gamma)(f(y)y^{-1})$ as desired.
Secondly, we compute the term $(\gamma\cdot f)(x)x^{-1}$ as
$(\gamma\cdot f)(x)x^{-1}$ $=(\gamma f\gamma^{-1})(x)x^{-1}$
$=\gamma(f(\gamma^{-1}(x))\gamma^{-1}(x^{-1})).$
By Lemma 3.7, the element $\beta\coloneqq x^{-\chi_{1}(\gamma^{-1})}\gamma^{-1}(x)$ is contained in the subgroup $\Pi(2,0)$. Since $f(\beta)\beta^{-1}$ is contained in $\Pi(\boldsymbol{m}+(2,0))$ by Lemma 3.5, we have
$\gamma(f(\gamma^{-1}(x))\gamma^{-1}(x^{-1}))$ $=\gamma(f(x^{\chi_{1}(\gamma^{-1})}\beta)\beta^{-1}x^{-\chi_{1}(\gamma^{-1})})$
$\equiv\gamma(f(x^{\chi_{1}(\gamma^{-1})})x^{-\chi_{1}(\gamma^{-1})})\bmod\Pi(\boldsymbol{m}+(2,0))$
$=\chi_{1}^{m_{1}+1}(\gamma)(f(x^{\chi_{1}(\gamma^{-1})})x^{-\chi_{1}(\gamma^{-1})})$
Hence, to obtain the desired equality, it suffices to show the congruence
$f(x^{n})x^{-n}\equiv(f(x)x^{-1})^{n}\bmod\Pi(\boldsymbol{m}+(2,0))$
for every $n\in\mathbb{Z}_{p}$. By continuity, it suffices to prove the congruence for every $n\in\mathbb{Z}_{\geq 1}$, which we prove by induction on $n$. We can compute the term $f(x^{n})x^{-n}$ as follows:
$f(x^{n})x^{-n}$ $=f(x)f(x^{n-1})x^{-(n-1)}x^{-1}$
$\equiv f(x)(f(x)x^{-1})^{n-1}x^{-1}\bmod\Pi(\boldsymbol{m}+(2,0))$
$=f(x)(f(x)x^{-1})^{n-1}f(x)^{-1}(f(x)x^{-1})=[f(x),(f(x)x^{-1})^{n-1}](f(x)x^{-1})^{n}$
$\equiv(f(x)x^{-1})^{n}\bmod\Pi(\boldsymbol{m}+(2,0)).$
Here, the induction hypothesis is used to deduce the second congruence, and we use
$[f(x),(f(x)x^{-1})^{n-1}]\in\Pi(\boldsymbol{m}+(2,0))$
to establish the last congruence. ∎
Remark 3.10.
Proposition 3.9 is an analogue of the fact that the map $\tilde{\iota}_{m}$ in [NT93, (1.11.2)] is compatible with the action of the product $\mathrm{GSp}(2g,n)\coloneqq\mathrm{GSp}(2g)\times\mathfrak{S}_{n}$ of the symplectic similitude group over $\mathbb{Z}_{p}$ and the $n$-th symmetric group, which is proved in [NT93, Theorem (1.11.4)].
By Proposition 3.9, we now understand how the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on $gr^{\boldsymbol{m}}_{1}\tilde{\Gamma}$. The following lemma allows us to describe the action of $\Gamma_{1}/F^{(1,0)}\Gamma$ on $gr^{\boldsymbol{m}}_{1}\Gamma$:
Lemma 3.11.
The natural homomorphism
$\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}\to\Gamma_{1}/F^{(1,0)}\Gamma$
is an isomorphism. Similarly, we have $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}\xrightarrow{\sim}\Gamma_{2}/F^{(0,1)}\Gamma$.
Proof.
Since $\ker(\Gamma^{{\dagger}}_{1}\to\Gamma_{1})=\ker(F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma)=\langle\mathrm{inn}(z)\rangle$, it suffices to prove that both $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$ and $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$ are both surjective by the snake lemma. First, we prove the surjective of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$. Let $\bar{\gamma}\in\Gamma_{1}$ be an arbitrary element and fix an arbitrary lift $\gamma\in\tilde{\Gamma}_{1}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, $g^{-1}\gamma g=\mathrm{inn}(g^{-1})\circ\gamma$ preserves $\langle z\rangle$. Moreover,
$(g^{-1}\gamma g)(y)y^{-1}=g^{-1}\gamma(y)gy^{-1}=[g^{-1},\gamma(y)]\gamma(y)y^{-1}\in\Pi(2).$
This shows $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$, hence the surjectivity of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$.
As for the surjectivity of $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$, let $\bar{\gamma}\in F^{(1,0)}\Gamma$ be arbitrary element and fix an arbitrary lift $\gamma\in F^{(1,0)}\tilde{\Gamma}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, then $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$ by the above argument. Observe that
$\gamma(z)z^{-1}=[g,z^{\alpha}]z^{\alpha-1}\in\Pi(3)\Rightarrow z^{\alpha-1}\in\Pi(3)\Rightarrow\alpha=1.$
Hence, to prove $(g^{-1}\gamma g)(x)x^{-1}=[g^{-1},\gamma(x)]\gamma(x)x^{-1}\in\Pi(2,0)$, it is enough to show $g\in\Pi(1,0)$. Note that $\Pi(\boldsymbol{1})/\Pi(2,1)=\Pi(\boldsymbol{1})/\Pi(2,0)\subset\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-submodule generated by $z=(y-1)x$. Hence $\gamma(z)z^{-1}=[g,z]=(g-1)z\in\Pi(2,0)$ implies $g=1$ in $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$, i.e. $g\in\Pi(1,0)$. This concludes the proof. ∎
To summarize the discussion so far, we obtain the following result on $\Gamma$:
Proposition 3.12.
For every $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, the action of $\Gamma_{1}/F^{(1,0)}\Gamma$ (resp. $\Gamma_{2}/F^{(0,1)}\Gamma$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
Proof.
The assertion follows from Proposition 3.9 and Lemma 3.11. ∎
The following lemma will be used in the next subsection:
Lemma 3.13.
We have
$F^{(1,0)}\Gamma^{{\dagger}}=\ker(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})\quad\text{and}\quad F^{(0,1)}\Gamma^{{\dagger}}=\ker(\Gamma^{{\dagger}}_{2}\xrightarrow{\chi_{2}}\mathbb{Z}_{p}^{\times}).$
Proof.
We only prove the first equality. Let $\gamma$ be an element of $\ker(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})$. Since we have $\gamma(y)y^{-1}\in\Pi(2)=\Pi(1,1)$ by definition of $\Gamma^{{\dagger}}_{1}$, it suffices to show $\gamma(x)=x\bmod\Pi(2,0)$.
Since the action of $\Gamma^{{\dagger}}_{1}$ on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by the image of $x$ is $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linear, the desired equality follows from the equality $\gamma(z)=z$. ∎
3.3. Two-variable filtration on Galois group
In the following, we apply various results obtained in this section to the case of the pro-$p$ fundamental group $\Pi=\Pi_{1,1}$. We use our fixed basis $x=x_{1}$ and $y=x_{2}$ of $\Pi_{1,1}$. Recall that we have the pro-$p$ outer Galois representation
$\rho_{X,p}\colon G_{K}\to\mathrm{Out}(\Pi_{1,1}),$
and the image of $\rho_{X,p}$ is contained in $\Gamma\subset\Gamma_{1,1}$ introduced in the last subsection.
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, let $F^{\boldsymbol{m}}G_{K}\subset F^{\lvert\boldsymbol{m}\rvert}G_{K}$ denote the inverse image of $F^{\boldsymbol{m}}\Gamma$ under $\rho_{X,p}$. We also define the associated graded quotients by
$\mathrm{gr}^{\boldsymbol{m}}_{1}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(1,0)}G_{K}\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(0,1)}G_{K}.$
Moreover, let $F_{1}G_{K}$ (resp. $F_{2}G_{K}$) denote the inverse image of $\Gamma_{1}$ (resp. $\Gamma_{2}$).
Lemma 3.14.
We have the following equalities:
  1. (1)
    $F_{1}G_{K}=G_{K(E[\bar{\mathfrak{p}}^{\infty}])}$ and $F_{2}G_{K}=G_{K(E[\mathfrak{p}^{\infty}])}$.
  2. (2)
    $F^{(1,0)}G_{K}=F^{(0,1)}G_{K}=F^{(1,1)}G_{K}=G_{K(E[p^{\infty}])}$.
Proof.
The assertion of (1) is clear. By Lemma 3.13 and Lemma 3.11, the group $F^{(1,0)}G_{K}$ coincides with the kernel of $F_{1}G_{K}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times}$, and thus it is equal to $G_{K(E[p^{\infty}])}$. The same argument shows $F^{(0,1)}G_{K}=G_{K(E[p^{\infty}])}$. Since we have
$F^{(1,0)}G_{K}\cap F^{(0,1)}G_{K}=F^{(1,1)}G_{K},$
the last assertion follows. ∎
Corollary 3.15.
For every $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, the action of $\mathrm{Gal}(K(E[p^{\infty}])/K(E[\bar{\mathfrak{p}}^{\infty}]))$ on $\mathrm{gr}^{\boldsymbol{m}}_{1}G_{K}$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\mathrm{Gal}(K(E[p^{\infty}])/K(E[\mathfrak{p}^{\infty}]))$ on $\mathrm{gr}^{\boldsymbol{m}}_{2}G_{K}$ is given by $\chi_{2}^{m_{2}}$.
Proof.
Both assertions immediately follow from Proposition 3.12 and Lemma 3.14. ∎
4. Proof of main theorem
In this section, we assume that $p$ satisfies the first two assumptions of Theorem 2.12, i.e. the class number of $K(p)$ is not divisible by $p$ and there are exactly two primes of $K(p^{\infty})$ above $p$. Moreover, we abbreviate $\mathrm{Gal}(\Omega/K(p))$ and $\mathrm{Gal}(\Omega^{\ast}/K(p))$ as $G$ and $G^{\ast}$, respectively.
Let us recall the definition of the index set
$I=\{\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 1}^{2}\setminus\{\boldsymbol{1}\}\mid m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert\}.$
In this section, we also use the following subset of $I$:
$I_{0}\coloneqq\{\boldsymbol{m}\in I\mid(p-1,p-1)\geq\boldsymbol{m}\}\cup\{(p,1),(1,p)\}.$
First, we give two filtrations on $G^{\ast}$. By Lemma 2.11, we have
$\rho_{X,p}(G_{K(p)})=\mathrm{Gal}(\Omega^{\ast}K(E[p])/K(p))=G^{\ast}\times\mathrm{Gal}(K(E[p])/K(p))\subset\Gamma_{1,1}.$
The image $\rho_{X,p}(G_{K(p)})$ has a descending central filtration defined at the beginning of Section 2.2. Moreover, it is also equipped with a two-variable filtration induced from the two-variable filtration on $\Gamma$ as in the previous subsection. By taking the images under the projection $\rho_{X,p}(G_{K(p)})\to G^{\ast}$, the group $G^{\ast}$ is also equipped with a descending central filtration $\{F^{m}G^{\ast}\}_{m\geq 1}$ and the two-variable filtration $\{F^{\boldsymbol{m}}G^{\ast}\}_{\boldsymbol{m}}$.
Moreover, by pulling back these filtrations under the projection $G\to G^{\ast}$, we obtain two filtrations $\{F^{m}G\}_{m\geq 1}$ and $\{F^{\boldsymbol{m}}G\}_{\boldsymbol{m}}$ on $G$. Then, by construction, the graded Lie algebras associated to $\{F^{m}G^{\ast}\}_{m\geq 1}$ and $\{F^{m}G\}_{m\geq 1}$ are naturally isomorphic to $\mathfrak{g}=\oplus_{m\geq 1}\mathfrak{g}_{m}$ defined in Section 2.2.
In the subsequent subsections, we give a proof of Theorem 2.12. The proof follows Sharifi’s approach [Sha02, Theorem 1.2]: We show that the intersection $\cap_{m\geq 1}F^{m}G=\mathrm{Gal}(\Omega/\Omega^{\ast})$ is trivial by the following strategy.
  1. (1)
    In Section 4.1, we construct an element $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G$ whose image in $\mathfrak{g}_{|\boldsymbol{m}|}$ satisfies the assumption of Conjecture 2.8. Moreover, from its construction, it is proved in Section 4.2 that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generates $F^{1}G$. Here, properties of two-variable filtrations are used in an essential way.
  2. (2)
    Conjecture 2.8 implies that the filtration $\{F^{m}G\}_{m\geq 1}$ coincides with the ‘‘fastest’’ descending central filtration $\{\tilde{F}^{m}G\}_{m\geq 1}$ satisfying $\sigma_{\boldsymbol{m}}\in\tilde{F}^{|\boldsymbol{m}|}G$ for every $\boldsymbol{m}\in I$. Since the intersection of the latter filtration is proved to be trivial in Lemma 4.11 (3), we obtain the desired result.
4.1. Construction of elements
In this subsection, we construct elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying the assumption of Conjecture 2.8 in such a way that can be regarded as a two-variable variant of Sharifi’s construction in [Sha02, 2]. We also construct auxiliary elements $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which are similar to $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ but more easier to handle group-theoretically. These auxiliary elements play an important role when proving that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generate $F^{1}G$ (cf. the proof of Theorem 2.12 in the next subsection)
First, we lift generators of $\mathrm{Gal}(K(p^{\infty})/K)\cong\Delta\times\mathbb{Z}_{p}^{2}$ to $\mathrm{Gal}(\Omega/K)$, as follows: Let us denote the maximal pro-$p$ subextension of $K(p^{\infty})/K$ by $K_{\infty}/K$. The upper exact sequence in the following commutative diagram
[…]
splits since $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a pro-$p$ group and $\Delta$ is a prime-to-$p$ group by Schur-Zassenhaus [Ser97, 5.9, Corollary 1]. We fix a section $r\colon\mathrm{Gal}(K(p)/K)\to\mathrm{Gal}(\Omega/K_{\infty})$ and identify $\mathrm{Gal}(\Omega/K)$ with the semi-direct product $G\rtimes\Delta$. Then the group $\Delta$ acts on $\mathrm{Gal}(\Omega/K(p^{\infty}))$ through this section by conjugation, and
$1\to\mathrm{Gal}(\Omega/K(p^{\infty}))\to G\to\mathrm{Gal}(K(p^{\infty})/K(p))\to 1$
is an exact sequence of pro-$p$ groups with $\Delta$-action, noting that $\Delta$ acts trivially on $\mathrm{Gal}(K(p^{\infty})/K(p))$. Then the sequence
$1\to\mathrm{Gal}(\Omega/K(p^{\infty}))^{\Delta}\to G^{\Delta}\to\mathrm{Gal}(K(p^{\infty})/K(p))\to 1$
obtained by taking $\Delta$-invariant subgroups is exact, since the homomorphism
$G^{\Delta}\to\mathrm{Gal}(K(p^{\infty})/K(p))$
is surjective by applying Lemma 4.1 in the case of $m=0$ below.
Let $\gamma_{1}$ (resp. $\gamma_{2}$) be an element of the $\Delta$-invariant subgroup $G^{\Delta}$ which restricts to a generator of $\mathrm{Gal}(K(p^{\infty})/K(\bar{\mathfrak{p}}^{\infty}\mathfrak{p}))\cong\mathbb{Z}_{p}$ (resp. $\mathrm{Gal}(K(p^{\infty})/K(\mathfrak{p}^{\infty}\bar{\mathfrak{p}}))\cong\mathbb{Z}_{p}$). Moreover, we fix a generator $\delta\in\mathbb{F}_{p}^{\times}$ and let $\delta_{1}$ (resp. $\delta_{2}$) be the image of $(\delta,1)$ (resp. $(1,\delta)$) in $\Delta$ through the isomorphism
$\Delta\xrightarrow[(\chi_{1},\chi_{2})]{\sim}(\mathbb{F}_{p}^{\times})^{\oplus 2}/\mathrm{im}(O_{K}^{\times})$
under the section $r\colon\Delta\to\mathrm{Gal}(\Omega/K)$. Then the element $\delta_{1}$ (resp. $\delta_{2}$) is contained in $\mathrm{Gal}(\Omega/K(\bar{\mathfrak{p}}))$ (resp. $\mathrm{Gal}(\Omega/K(\bar{p}))$) and, by construction, we have the following relations:
$[\delta_{1},\delta_{2}]=1,\quad[\delta_{1},\gamma_{1}]=1,\quad[\delta_{1},\gamma_{2}]=1,\quad[\delta_{2},\gamma_{1}]=1\quad\text{and}\quad[\delta_{2},\gamma_{2}]=1.\quad$
For $m\in\mathbb{Z}_{\geq 0}$ and $i=1,2$, we define $\epsilon_{i,m}\in\mathbb{Z}_{p}[\mathrm{Gal}(\Omega/K)]$ to be
$\epsilon_{i,m}\coloneqq\frac{1}{p-1}\sum_{j=0}^{p-2}\chi_{i}^{-m}(\delta_{i}^{j})\delta_{i}^{j}.$
Moreover, for $g\in G$, we define
$g^{\epsilon_{i,m}}\coloneqq(g\cdot\delta_{i}g^{\chi_{i}^{-m}(\delta_{i})}\delta_{i}^{-1}\cdots\delta_{i}^{p-2}g^{\chi_{i}^{-m}(\delta_{i}^{p-2})}\delta_{i}^{-(p-2)})^{\frac{1}{p-1}},$
and let $g^{\epsilon_{i,m}^{j}}=(\cdots(g^{\epsilon_{i,m}})\cdots)^{\epsilon_{i,m}}$ denote its $j$-th iterate for $j\geq 1$. Since $G$ is nonabelian, we do not necessarily have the equality $g^{\epsilon_{i,m}^{j}}=g^{\epsilon_{i,m}}$. However, we have the following lemma, originally due to Sharifi, which is one of the key ingredients to construct desired elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ group-theoretically.
Lemma 4.1 (cf. Sharifi [Sha02, Lemma 2.1]).
For every $g\in G$, $m\in\mathbb{Z}_{\geq 0}$ and $i=1,2$, the limit
$g^{(i,m)}\coloneqq\lim_{j\to\infty}g^{\epsilon_{i,m}^{j}}$
exists and satisfies $\delta_{i}g^{(i,m)}\delta_{i}^{-1}\coloneqq(g^{(i,m)})^{\chi_{i}^{m}(\delta_{i})}.$
Proof.
The assertion follows from the same argument as [Sha02, Lemma 2.1]. ∎
Let $A$ denote the maximal abelian quotient $\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$ of $F^{1}G=\mathrm{Gal}(\Omega/K(p^{\infty}))$. Then $A$ is naturally endowed with the structure of a $\Lambda\coloneqq\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$-module. In the following lemma and remark, we determine the structure of this $\Lambda$-module.
Lemma 4.2.
Assume that the class number of $K(p)$ is not divisible by $p$, and there are only two primes of $K(p^{\infty})$ above $p$. Let $\boldsymbol{m}\in I_{0}$ and $A^{\boldsymbol{m}}\coloneqq\epsilon_{1,m_{1}}\epsilon_{2,m_{2}}A$. Then the following assertions hold:
  1. (1)
    If $\boldsymbol{m}\neq(1,p)$ or $(p,1)$, then $A^{\boldsymbol{m}}$ is a cyclic $\Lambda$-module.
  2. (2)
    $A^{(p,1)}=A^{(1,p)}$ is a quotient of the annihilator ideal $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ of the Tate twist $\mathbb{Z}_{p}(1)$, i.e. a quotient of the $\Lambda$-module $M$ generated by two elements $v_{1},v_{2}$ satisfying $(\gamma_{1}-\chi_{1}(\gamma_{1}))v_{2}=(\gamma_{2}-\chi_{2}(\gamma_{2}))v_{1}$.
Proof.
We have the five-term exact sequence between $\mathbb{F}_{p}[\Delta]$-modules [NSW08, (1.6.7)]
00 0 $\to H^{1}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{1}(G,\mathbb{F}_{p})\to\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})$
$\to H^{2}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{2}(G,\mathbb{F}_{p}).$
Here we note that
$H^{1}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{F}_{p})^{\mathrm{Gal}(K(p^{\infty})/K(p))}$ $=\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})$
$=\oplus_{\boldsymbol{m}\in I_{0}\setminus\{(p,1)\}}\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A^{\boldsymbol{m}},\mathbb{F}_{p}).$
To prove the assertion of (1), it suffices to compute the dimension of each eigenspace of this cohomology group. First, we compute the dimension of each eigenspace of $H^{1}(G,\mathbb{F}_{p})$. The Kummer exact sequence $1\to\mu_{p}\to\mathbb{G}_{m}\to\mathbb{G}_{m}\to 1$ gives
$1\to O_{K(p)}[1/p]^{\times}/p\to H^{1}(G,\mu_{p})\to H^{1}_{\mathrm{\acute{e}t}}(O_{K(p)}[1/p],\mathbb{G}_{m})[p]\to 0$
Since the class number of $K(p)$ is prime to $p$, the group $H^{1}_{\mathrm{\acute{e}t}}(O_{K(p)}[1/p],\mathbb{G}_{m})[p]$, which is nothing but the $p$-torsion subgroup of the $p$-class group, is trivial. Hence we obtain an isomorphism
$O_{K(p)}[1/p]^{\times}/p\xrightarrow{\sim}H^{1}(G,\mu_{p}).$
[…]
$O_{K(p)}[1/p]^{\times}\otimes\mathbb{Q}\cong\mathbb{Q}[\Delta]\oplus\mathbb{Q}.$
By decomposing the $p$-unit group into the product of the torsion-part and the free-part, it follows that the dimension of $\chi^{\boldsymbol{m}}$-component of $H^{1}(G,\mathbb{F}_{p})$ is at most one if $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$ and otherwise at most two. By counting argument, all these inequalities are equalities. Since the action of $\Delta$ on $H^{i}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})$ ($i=1,2$) is trivial, (1) follows from the five-term exact sequence.
By our assumption, there exists a unique prime of $K(p^{\infty})$ above $\mathfrak{p}$ which we also denote by the same letter $\mathfrak{p}$. In other words, we have an isomorphism
$\mathrm{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})\cong\mathrm{Gal}(K(p^{\infty})/K(p)).$
Moreover, by a result of Wintenberger [Win81, THÉORÈME], the decomposition group of $A^{(1,p)}$ at $\mathfrak{p}$ is isomorphic to
$\mathrm{ann}_{\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})]]}(\mathbb{Z}_{p}(1))=\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1)).$
Hence to prove (2), it suffices to show that the decomposition group of $A^{(1,p)}$ at $\mathfrak{p}$ coincides with the whole $A^{(1,p)}$. First, note that
$\mathrm{Hom}_{\Delta}(A^{(1,p)},\mathbb{F}_{p}(1))\cong\mathrm{Hom}_{\Delta}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p}(1))\cong O_{K}[1/p]^{\times}/p$
is generated by images of $\pi$ and $\bar{\pi}$. Hence the assertion is equivalent to saying that the images of $\pi$ and $\bar{\pi}$ in $K_{\mathfrak{p}}^{\times}/(K_{\mathfrak{p}}^{\times})^{p}$ is still two-dimensional. To prove the latter assertion, it suffices to show that the projection of $\bar{\pi}\in O_{K_{\mathfrak{p}}}^{\times}$ to the group of principal units $1+\pi O_{K_{\mathfrak{p}}}\cong\mathbb{Z}_{p}$ is a generator. Since the Frobenius element $\mathrm{Frob}_{\bar{\mathfrak{p}}}\in\mathrm{Gal}(K(\mathfrak{p}^{\infty})/K)\cong O_{K_{\mathfrak{p}}}^{\times}/O_{K}^{\times}$ at $\bar{\mathfrak{p}}$ is a generator by our assumption on the number of primes above $p$, we obtain the desired assertion. ∎
In fact, we can completely determine the $\Lambda$-module structure of $A$ and the structure of the pro-$p$ group $\mathrm{Gal}(\Omega/K(p^{\infty}))$, though they are not necessary to prove Theorem 2.12:
Remark 4.3.
(1) We keep the same assumption as Lemma 4.2. It follows that there is a surjection
$f\colon\Lambda^{[K(p):K]-1}\oplus\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\to\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$
between $\Lambda$-modules. We claim that $f$ is an isomorphism. Let $\mathrm{Frac}(\Lambda)$ denote the field of fractions of $\Lambda$. Then we have
$\dim_{\mathrm{Frac}(\Omega)}\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\otimes_{\Lambda}\mathrm{Frac}(\Omega)=1.$
Moreover, by [NQD84, Corollaire 2.7], we have
$\dim_{\mathrm{Frac}(\Omega)}\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}\otimes_{\Lambda}\mathrm{Frac}(\Omega)=[K(p):K].$
Hence it follows that $\ker(f)$ is a torsion $\Lambda$-module. However, the $\Lambda$-module $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ is torsion-free since it is an ideal of $\Lambda$. Hence the kernel of $f$ is trivial, as desired.
(2) As a consequence of the previous remark, the $\Lambda$-module $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ is $\mathbb{Z}_{p}$-torsion free. Moreover, we have $H^{2}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{Q}_{p}/\mathbb{Z}_{p})=0$, i.e. the weak Leopoldt conjecture holds for the $\mathbb{Z}_{p}^{2}$-extension $K(p^{\infty})/K(p)$, since it contains the cyclotomic $\mathbb{Z}_{p}$-extension by [NQD84, Theoreme 2.2]. Now, by considering a long exact sequence associated to
$0\to\mathbb{F}_{p}\to\mathbb{Q}_{p}/\mathbb{Z}_{p}\xrightarrow{\times p}\mathbb{Q}_{p}/\mathbb{Z}_{p}\to 0,$
it follows that $H^{2}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{F}_{p})=0$. In other words, the pro-$p$ group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a free pro-$p$ group. The proof of Theorem 2.12 below will provide an explicit basis of this free pro-$p$ group.
Now we define elements $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G_{K}$ for $\boldsymbol{m}\in I_{0}$ as follows:
Construction. For $\boldsymbol{m}\in I_{0}$, we choose an element $t_{\boldsymbol{m}}\in\mathrm{Gal}(\Omega/K(p^{\infty}))$, as follows:
  • If $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$, we choose a lift $t_{\boldsymbol{m}}\in\mathrm{Gal}(\Omega/K(p^{\infty}))$ of a generator of $A^{\boldsymbol{m}}$ as $\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$-module.
  • For $\boldsymbol{m}=(p,1)$ and $(1,p)$, fix a surjection from $M$ given above to $A^{(p,1)}$ and let $t_{(p,1)}$ and $t_{(1,p)}$ be arbitrary lifts of the images of $v_{1}$ and $v_{2}$, respectively.
  • For $\boldsymbol{m}=(p-1,p-1)$, set $t_{(p-1,p-1)}=[\gamma_{1},\gamma_{2}]$.
For every $\boldsymbol{m}\in I_{0}$, let
$\sigma_{\boldsymbol{m}}\coloneqq\left(t_{\boldsymbol{m}}^{(1,m_{1})}\right)^{(2,m_{2})}\quad\text{and}\quad g_{\boldsymbol{m}}\coloneqq\sigma_{\boldsymbol{m}}.$
Note that $\sigma_{(p-1,p-1)}=g_{(p-1,p-1)}=[\gamma_{1},\gamma_{2}]$ since $\gamma_{1}$ and $\gamma_{2}$ commute with $\delta_{1}$ and $\delta_{2}$. The elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$ satisfy the following properties:
Lemma 4.4.
For $\boldsymbol{m}\in I_{0}$, the following two assertions hold.
  1. (1)
    The element $\sigma_{\boldsymbol{m}}$ is contained in $F^{\boldsymbol{m}}G$, and its image in $\mathfrak{g}_{\lvert\boldsymbol{m}\rvert}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component.
  2. (2)
    The element $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})$ generates $\kappa_{\boldsymbol{m}}(F^{1}G)$, which is nonzero.
Before proving the lemma, we first prove the following two lemmas concerning the case where $m_{1}=1$ or $m_{2}=1$.
Lemma 4.5.
Let $n\geq 2$ be an arbitrary integer such that $n\equiv 1\bmod p-1$. Then we have $\kappa_{(1,n)}(t_{(p,1)})=0$. Similarly, for such $n$, we have $\kappa_{(n,1)}(t_{(1,p)})=0$.
Proof.
Note that $\kappa_{(n,1)}$ factors through $A^{(1,1)}$, and we have
$(\gamma_{2}-\chi_{2}(\gamma_{2}))t_{(p,1)}=(\gamma_{1}-\chi_{1}(\gamma_{1}))t_{(1,p)}$
in $A^{(1,1)}$ by definitions of $t_{(p,1)}$ and $t_{(1,p)}$. Hence we have
$\kappa_{(1,n)}((\gamma_{2}-\chi_{2}(\gamma_{2}))t_{(p,1)})$ $=(\chi_{2}^{n}(\gamma_{2})-\chi_{2}(\gamma_{2}))\kappa_{(1,n)}(t_{(p,1)})$
$=\kappa_{(1,n)}((\gamma_{1}-\chi_{1}(\gamma_{1}))t_{(1,p)})$
$=(\chi_{1}(\gamma_{1})-\chi_{1}(\gamma_{1}))\kappa_{p,1}(t_{(1,p)})=0.$
Since $\chi_{2}^{n-1}(\gamma_{2})\neq 1$, it follows that $\kappa_{(1,n)}(t_{(p,1)})=0$. The same argument shows $\kappa_{(n,1)}(t_{(1,p)})=0$, as desired. ∎
Lemma 4.6.
We have $t_{(p,1)}\in F^{(2,1)}G$ and $t_{(1,p)}\in F^{(1,2)}G$.
Proof.
Let $s\colon F^{1}G_{K}\to F^{3}\Gamma^{{\dagger}}_{1,1}$ be the lift of $\rho_{X,p}$ constructed in Section 2.1. Then the map $s$ factors through $F^{1}G_{K}\to F^{1}G$, and we denote the resulting homomorphism $F^{1}G\to F^{3}\Gamma^{{\dagger}}_{1,1}$ by the same letter. By Lemma 3.4, to show $t_{(p,1)}\in F^{(2,1)}G$, it suffices to prove that $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$ and $s(t_{(p,1)})(x_{1})x_{1}^{-1}\in\Pi(3,1)$.
First, we prove that $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$. To verify this inclusion, it is enough to show that the power series $\alpha_{1,1}(t_{(p,1)})$ is divisible by $T_{1}$. In fact, then we would have
$s(t_{(p,1)})(x_{2})x_{2}^{-1}$ $=\alpha_{1,1}(t_{(p,1)})[x_{2},z]$
$=\frac{\alpha_{1,1}(t_{(p,1)})}{T_{1}}[x_{1},[x_{2},z]]\in\Pi(2)/[\Pi(2),\Pi(2)].$
by definition of $\alpha_{1,1}$. Since it holds that
$[\Pi(2),\Pi(2)]=[\Pi(1,1),\Pi(1,1)]\subset\Pi(2,2)$
and $[x_{1},[x_{2},z]]\in\Pi(2,2)$, we have $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$ as desired.
In view of the explicit formula for $\alpha_{1,1}$ (Theorem 2.3), it suffices to show that
$\kappa_{\boldsymbol{n}}(t_{(p,1)})=0\quad\text{for every }\boldsymbol{n}=(1,n_{2})\in I.$
For $\boldsymbol{n}=(1,n_{2})\in I$, we have $\kappa_{\boldsymbol{n}}(t_{(p,1)})=0$ unless $n_{2}\equiv 1\bmod p-1$, since $\kappa_{(1,n)}$ factors through $A^{(1,n)}$ while $t_{(p,1)}$ is a lift of an element of $A^{(p,1)}=A^{(1,p)}$. Moreover, by Lemma 4.5 above, we also have $\kappa_{\boldsymbol{n}}(t_{(p,1)})=0$ when $n_{2}\equiv 1\bmod p-1$. This implies that $\alpha_{1,1}(t_{(p,1)})$ is divisible by $T_{1}$ as desired.
In the following, we prove $s(t_{(p,1)})(x_{1})x_{1}^{-1}\in\Pi(3,1)=\Pi(3,0)$ (here, we use Lemma 3.2). We have already shown that this element is contained in $\Pi(2,0)$ and, from the equality $s(t_{(p,1)})(z)=z$ and $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$, it follows that
$[s(t_{(p,1)})(x_{1})x_{1}^{-1},x_{2}]\in\Pi(3,2)\subset\Pi(3,0).$
In particular, we have the following:
$s(t_{(p,1)})(x_{1})x_{1}^{-1}\in\Pi(2,0)/\Pi(3,0)$ is invariant under the conjugation of $x_{2}$.
Hence it suffices to show the following claim:
Claim. The identity is the only element of $\Pi(2,0)/\Pi(3,0)$ which is invariant under the conjugation of $x_{2}$.
To prove the claim, first note that, for every $m\geq 1$, the group $\Pi(m,0)$ is just the $m$-th component of the descending central series of $\Pi(1,0)$, which is a free pro-$p$ group on the set $\{w_{n}\}_{n\geq 1}$ where $w_{0}\coloneqq x_{1}$ and $w_{n}\coloneqq[x_{2},w_{n-1}]$ for every $n\geq 1$.
[…]
$F_{n}/F_{n}(2)\wedge F_{n}/F_{n}(2)\xrightarrow{\sim}F_{n}(2)/F_{n}(3)$
for every $n\geq 1$, since the Lie algebra associated to the descending central series of $F_{n}$ is freely generated by the image of $\{w_{i}\}_{0\leq i<n}$ in $F_{n}/F_{n}(2)$. In other words, the quotient $F_{n}(2)/F_{n}(3)$ is a free $\mathbb{Z}_{p}$-module with basis $\{[w_{i},w_{j}]\}_{0\leq i<j<n}$. Therefore, we have
[…]
Observe that the action of $x_{2}$ sends $[w_{i},w_{j}]\in\Pi(2,0)/\Pi(3,0)$ to
$[w_{i+1}w_{i},w_{j+1}w_{j}]=[w_{i+1},w_{j+1}]+[w_{i+1},w_{j}]+[w_{i},w_{j+1}]+[w_{i},w_{j}].$
If $v=(v_{i,j})\in\prod_{0\leq i<j}\mathbb{Z}_{p}[w_{i},w_{j}]=\Pi(2,0)/\Pi(3,0)$ is invariant under the conjugation of $x_{2}$, then one can show that $v_{0,j}=0$ for every $j>0$ by induction on $j$. By repeating induction for every $i>0$, it follows that $v_{i,j}=0$ for every $0\leq i<j$, hence $v=0$ as desired. ∎
Proof of Lemma 4.4.
First, note that $\sigma_{\boldsymbol{m}}\in F^{1}G=F^{\boldsymbol{1}}G$ by Lemma 3.14. Hence the assertion of (1) immediately follows from Corollary 3.15, except when $\boldsymbol{m}=(1,p)$ or $(p,1)$. If $\boldsymbol{m}=(p,1)$, we know $t_{(p,1)}\in F^{(2,1)}G$ by Lemma 4.6. Then the claim that $\sigma_{(p,1)}\in F^{(p,1)}G$ follows from Corollary 3.15, and the case where $\boldsymbol{m}=(1,p)$ is similar. The second assertion of (1) also follows from Corollary 3.15.
Next we prove the assertion of (2). For $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$, the assertion immediately follows since we have $\kappa_{\boldsymbol{m}}(t_{\boldsymbol{m}})=\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})$ and the element $t_{\boldsymbol{m}}$ generates $A^{\boldsymbol{m}}$ as a $\Lambda$-module. Since the $\boldsymbol{m}$-th Soulé character is nontrivial by Theorem 2.6 (2), the image $\kappa_{\boldsymbol{m}}(F^{1}G)$ is nonzero.
Assume that $\boldsymbol{m}=(p-1,p-1)$. Since $A^{(p-1,p-1)}=A^{\Delta}$, the $\Delta$-invariant part of $A$, it follows that $A^{(p-1,p-1)}$ is isomorphic to $\mathrm{Gal}(L/K_{\infty})^{{\rm ab}}$, where $L$ is the maximal pro-$p$ extension of $K$ unramified outside $p$. Since $K$ has class number one, it follows by [NSW08, (10.7.13) Theorem] that $\mathrm{Gal}(L/K)$ is a free pro-$p$ group of rank two on the set $\{\gamma_{1},\gamma_{2}\}$. Hence $\mathrm{Gal}(L/K_{\infty})^{\rm{ab}}$, which is just the maximal abelian quotient of the commutator subgroup of a free pro-$p$ group of rank two, is a free $\Lambda$-module of rank one generated by $[\gamma_{1},\gamma_{2}]$ by [Iha86, Theorem 2]. Hence the image of $\kappa_{(p-1,p-1)}(\sigma_{(p-1,p-1)})$ generates $\kappa_{(p-1,p-1)}(A)$, which is nonzero by Theorem 2.6 (2). Finally, the case where $\boldsymbol{m}=(p,1)$ or $(1,p)$ follows from Lemma 4.5 and Theorem 2.6 (2). ∎
We inductively define $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ for general $\boldsymbol{m}\in I$ as follows:
Construction.
  • First, assume that $\boldsymbol{m}=(m_{1},m_{2})\in I$ is an index satisfying that $m_{1}\geq p$, $m_{2}\leq p-1$ and $\boldsymbol{m}\not\equiv\boldsymbol{1}\bmod p-1$. we define $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ as
    $\sigma_{\boldsymbol{m}}\coloneqq\left(\gamma_{1}\sigma_{\boldsymbol{m}-(p-1,0)}\gamma_{1}^{-1}\sigma_{\boldsymbol{m}-(p-1,0)}^{-\chi_{1}^{m_{1}}(\gamma_{1})}\right)^{(1,m_{1})}$
    and
    $g_{\boldsymbol{m}}\coloneqq\gamma_{1}g_{\boldsymbol{m}-(p-1,0)}\gamma_{1}^{-1}g_{\boldsymbol{m}-(p-1,0)}^{-\chi_{1}^{m_{1}}(\gamma_{1})}.$
  • Secondly, if $\boldsymbol{m}\in I$ is an index such that $m_{2}\geq p$ and $\boldsymbol{m}\not\equiv\boldsymbol{1}\bmod p-1$, we define $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ as
    $\sigma_{\boldsymbol{m}}\coloneqq\left(\gamma_{2}\sigma_{\boldsymbol{m}-(0,p-1)}\gamma_{2}^{-1}\sigma_{\boldsymbol{m}-(0,p-1)}^{-\chi_{2}^{m_{2}}(\gamma_{2})}\right)^{(2,m_{2})}$
    and
    $g_{\boldsymbol{m}}\coloneqq\gamma_{2}g_{\boldsymbol{m}-(0,p-1)}\gamma_{2}^{-1}g_{\boldsymbol{m}-(0,p-1)}^{-\chi_{2}^{m_{2}}(\gamma_{2})}.$
Finally, we consider the case where $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$. In this case, if $\boldsymbol{m}\geq(2,2)$, we obtain two candidates for $\sigma_{\boldsymbol{m}}$ obtained from applying the construction inductively from $\sigma_{(1,p)}$ and $\sigma_{(p,1)}$.
However, these candidates define the same element on the level of the abelianization $A$ of the Galois group of $\Omega$ over $K(p^{\infty})$. In fact, write $\boldsymbol{m}=(1+n_{1}(p-1),1+n_{2}(p-1))$ for some $\boldsymbol{n}=(n_{1},n_{2})\in\mathbb{Z}_{\geq 1}^{2}$. If we start from $\sigma_{(p,1)}$ to obtain $\sigma_{\boldsymbol{m}}$ through the above construction, we have
$\sigma_{\boldsymbol{m}}=\prod_{i=1}^{n_{1}-1}(\gamma_{1}-\chi_{1}(\gamma_{1})^{1+i(p-1)})\prod_{j=0}^{n_{2}-1}(\gamma_{2}-\chi_{2}(\gamma_{2})^{1+j(p-1)})\sigma_{(p,1)}.$
as an element of $A$. On the other hand, if we start from $\sigma_{(p,1)}$ we have
$\sigma_{\boldsymbol{m}}=\prod_{i=0}^{n_{1}-1}(\gamma_{1}-\chi_{1}(\gamma_{1})^{1+i(p-1)})\prod_{j=1}^{n_{2}-1}(\gamma_{2}-\chi_{2}(\gamma_{2})^{1+j(p-1)})\sigma_{(1,p)}.$
Hence two candidates yield the same element on $A$ for every $\boldsymbol{n}\geq\boldsymbol{1}$ by Lemma 4.2(2), and a similar argument can also be applied to the case of $g_{\boldsymbol{m}}$. Now we define $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ for $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$ as follows.
  • For an index $\boldsymbol{m}=(m,1)$ such that $m\geq 2$ and $m\equiv 1\bmod p-1$, we define $\sigma_{(m,1)}$ and $g_{(m,1)}$ by applying the above construction, starting from $\sigma_{(p,1)}$. Similarly, for every index $\boldsymbol{m}=(1,m)$ such that $m\geq 2$ and $m\equiv 1\bmod p-1$, we define $\sigma_{(1,m)}$ and $g_{(1,m)}$ by applying the above construction, starting from $\sigma_{(1,p)}$.
  • Let $\boldsymbol{m}\in I$ be an index satisfying $\boldsymbol{m}\geq(2,2)$ and $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$. Then we define $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ by applying the above construction, starting from $\sigma_{(p,1)}$.
We have the following lemma.
Lemma 4.7.
For $\boldsymbol{m}\in I$, two elements $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ define the same element of $A$.
Proof.
Since $A$ is abelian, the elements $\epsilon_{1,m}$ and $\epsilon_{2,m}$ act on $A$ as idempotents for every $m\geq 0$. Moreover, the action of $\epsilon_{i,m}$ commutes with the conjugation by $\gamma_{1}$ and $\gamma_{2}$. Hence the assertion follows from the construction of $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$. ∎
We now show that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfy the assumption of Conjecture 2.8:
Proposition 4.8.
For $\boldsymbol{m}=(m_{1},m_{2})\in I$, the following assertions hold.
  1. (1)
    The element $\sigma_{\boldsymbol{m}}$ is contained in $F^{\boldsymbol{m}}G$, and the image of $\sigma_{\boldsymbol{m}}$ in the $|\boldsymbol{m}|$-th graded quotient $\mathfrak{g}_{\lvert\boldsymbol{m}\rvert}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component of $\mathfrak{g}_{\lvert\boldsymbol{m}\rvert}$.
  2. (2)
    The element $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})$ generates $\kappa_{\boldsymbol{m}}(F^{|\boldsymbol{m}|}G)$, which is nonzero.
Proof.
First, by Lemma 4.4, the assertion of (1) holds for every $\boldsymbol{m}\in I_{0}$. For every $\boldsymbol{m}\in I$ satisfying the assertion of (1), we have $\gamma_{1}\sigma_{\boldsymbol{m}}\gamma_{1}\sigma_{\boldsymbol{m}}^{-\chi_{1}^{m_{1}}(\gamma_{1})}\in F^{\boldsymbol{m}+(1,0)}G$, by Corollary 3.15. Now we claim that
$\sigma_{\boldsymbol{m}+(p-1,0)}=\left(\gamma_{1}\sigma_{\boldsymbol{m}}\gamma_{1}\sigma_{\boldsymbol{m}}^{-\chi_{1}^{m_{1}}(\gamma_{1})}\right)^{(1,m_{1})}\in F^{\boldsymbol{m}+(p-1,0)}G.$
The element $\sigma_{\boldsymbol{m}+(p-1,0)}$ is contained in $F^{\boldsymbol{m}+(1,0)}G$. By the construction of $\sigma_{\boldsymbol{m}+(p-1,0)}$ (cf. Lemma 4.1), we have the equality
$\delta_{1}\sigma_{\boldsymbol{m}+(p-1,0)}\delta_{1}^{-1}=\chi_{1}^{m_{1}}(\delta_{1})\sigma_{\boldsymbol{m}+(p-1,0)}$
in the two-variable graded quotient $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}G$. However, by Corollary 3.15, we also have
$\delta_{1}\sigma_{\boldsymbol{m}+(p-1,0)}\delta_{1}^{-1}=\chi_{1}^{m_{1}+1}(\delta_{1})\sigma_{\boldsymbol{m}+(p-1,0)}$
in the same graded quotient. Since the element $\chi_{1}(\delta_{1})\in\mathbb{Z}_{p}^{\times}$ is of order $p-1$, we have $\sigma_{\boldsymbol{m}+(p-1,0)}\in F^{\boldsymbol{m}+(2,0)}G$. By applying the same argument to graded quotients
$\mathrm{gr}^{\boldsymbol{m}+(2,0)}_{1}G,\dots,\mathrm{gr}^{\boldsymbol{m}+(p-2,0)}_{1}G,$
we obtain $\sigma_{\boldsymbol{m}+(p-1,0)}\in F^{\boldsymbol{m}+(p-1,0)}G$, as desired. A similar argument immediately shows that $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G$ for every $\boldsymbol{m}\in I$. This proves the former assertion of (1), and the latter assertion follows from Corollary 3.15.
To prove the assertion of (2), it suffices to show that the element $\kappa_{\boldsymbol{m}}(g_{\boldsymbol{m}})$ generates $\kappa_{\boldsymbol{m}}(F^{|\boldsymbol{m}|}G)$ for every $\boldsymbol{m}\in I$ by Lemma 4.7. We fix an arbitrary index $\boldsymbol{m}_{0}\in I_{0}\setminus\{(1,p)\}$, and prove the assertion of (2) for every $\boldsymbol{m}$ such that $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$. By Lemma 4.2, the $\Lambda$-module $A^{\boldsymbol{m}_{0}}$ is generated by (the image of) $g_{\boldsymbol{m}_{0}}$ if $\boldsymbol{m}_{0}\neq(p,1)$, and generated by $g_{(p,1)}$ and $g_{(1,p)}$ if $\boldsymbol{m}_{0}=(p,1)$.
For every $m\geq\lvert\boldsymbol{m}_{0}\rvert$ such that $m\equiv\lvert\boldsymbol{m}_{0}\rvert\bmod p-1$, let $A^{\boldsymbol{m}_{0},m}$ be the image of $F^{m}G$ in $A^{\boldsymbol{m}_{0}}$. Note the image of $F^{m+1}G$ in $A^{\boldsymbol{m}_{0},m}$ coincides with $A^{\boldsymbol{m}_{0},m+(p-1)}$. In fact, assume that there is an integer $11\leq i<p-$ 1 such that the image of $F^{m+i}G$ properly contains $A^{\boldsymbol{m}_{0},m+(p-1)}$, and take $i$ to be maximal among such integers. Then we would have a nontrivial homomorphism
$\mathfrak{g}_{m+i}\to A^{\boldsymbol{m}_{0},m}/A^{\boldsymbol{m}_{0},m+p-1}.$
However, since $\mathfrak{g}_{m+i}$ is embedded in $\mathfrak{g}_{m+i}\otimes\mathbb{Q}_{p}$, which is a direct sum of $\mathbb{Q}_{p}(n_{1},n_{2})$ where $n_{1}+n_{2}=m+i\not\equiv m\bmod p-1$, the subspace $\epsilon_{1,m_{1}}\epsilon_{2,m_{2}}\mathfrak{g}_{m+i}$ is trivial. Hence the image of the above homomorphism should be trivial, which is a contradiction.
By what we have just proved, we have a homomorphism
$\mathfrak{g}_{|\boldsymbol{m}|}\to A^{\boldsymbol{m}_{0},|\boldsymbol{m}|}/A^{\boldsymbol{m}_{0},|\boldsymbol{m}|+p-1}$
sending $\sigma_{\boldsymbol{m}}$ to $g_{\boldsymbol{m}}$, and the character $\kappa_{\boldsymbol{m}}\mid_{F^{\lvert\boldsymbol{m}\rvert}G}$ factors through $A^{\boldsymbol{m}_{0},\lvert\boldsymbol{m}\rvert}/A^{\boldsymbol{m}_{0},\lvert\boldsymbol{m}\rvert+(p-1)}$, for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$. Now we prove the following claim:
Claim. The $\Lambda$-module $A^{\boldsymbol{m}_{0},m}$ is generated by $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}}$, where $\boldsymbol{m}$ ranges over the indexes $\boldsymbol{m}\in I$ such that $|\boldsymbol{m}|=m$ and $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$. Moreover, the element $\kappa_{\boldsymbol{m}}(g_{\boldsymbol{m}})$ generates the image of $\kappa_{\boldsymbol{m}}(A^{\boldsymbol{m}_{0},m})\neq 0$.
Before giving the proof of the claim, we remark that, for every $\boldsymbol{m}\in I$ with $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$, the image of $g_{\boldsymbol{m}}$ in $A^{\boldsymbol{m}_{0},\lvert\boldsymbol{m}\rvert}/A^{\boldsymbol{m}_{0},\lvert\boldsymbol{m}\rvert+(p-1)}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component, since it is the image of $\sigma_{\boldsymbol{m}}\in\mathfrak{g}_{|\boldsymbol{m}|}$, which is contained in the $\chi^{\boldsymbol{m}}$-isotypic component by (1). In particular, the second half of the claim follows from the first.
In the following, we prove the first half of the claim by induction on $m$. We identify the ring $\Lambda=\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$ with the power series ring $\mathbb{Z}_{p}[[S_{1},S_{2}]]$ via $S_{i}=\gamma_{i}-1$ for $i=1,2$.
First, we know that the claim holds for $m=|\boldsymbol{m}_{0}|$. We prove the claim for $m+(p-1)$, assuming it is true for $m$. Take an arbitrary element $x\in A^{\boldsymbol{m}_{0},m+(p-1)}$. By induction hypothesis, the element $x$ can be written as
$x=\sum f_{\boldsymbol{m}}(S_{1},S_{2})g_{\boldsymbol{m}}$
for some $f_{\boldsymbol{m}}(S_{1},S_{2})\in\mathbb{Z}_{p}[[S_{1},S_{2}]]$ where $\boldsymbol{m}$ ranges over the indexes $\boldsymbol{m}\in I$ satisfying $|\boldsymbol{m}|=m$ and $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$.
Since $x\in A^{\boldsymbol{m}_{0},m+(p-1)}$, we have $\kappa_{\boldsymbol{n}}(x)=0$ for every $\boldsymbol{n}=(n_{1},n_{2})$ such that $|\boldsymbol{n}|=m$ and $\boldsymbol{n}\equiv\boldsymbol{m}_{0}\bmod p-1$, by Lemma 2.7 (2). Hence we have
$\kappa_{\boldsymbol{n}}(x)$ $=\sum\kappa_{\boldsymbol{n}}(f_{\boldsymbol{m}}(S_{1},S_{2})g_{\boldsymbol{m}})$
$=\sum f_{\boldsymbol{m}}(\chi_{1}^{n_{1}}(\gamma_{1})-1,\chi_{2}^{n_{2}}(\gamma_{2})-1)\kappa_{\boldsymbol{n}}(g_{\boldsymbol{m}})$
$=f_{\boldsymbol{n}}(\chi_{1}^{n_{1}}(\gamma_{1})-1,\chi_{2}^{n_{2}}(\gamma_{2})-1)\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})=0.$
Note that $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})$ generates $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)$ by induction hypothesis. Moreover, the submodule $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)\subset\mathbb{Z}_{p}(\boldsymbol{n})$ is nonzero by Theorem 2.6 (2) and Corollary A.3. Therefore we have $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})\neq 0$, implying that $f_{\boldsymbol{n}}(\chi_{1}^{n_{1}}(\gamma_{1})-1,\chi_{2}^{n_{2}}(\gamma_{2})-1)=0$. Equivalently, the power series $f_{\boldsymbol{n}}(S_{1},S_{2})$ is contained in the ideal $(S_{1}-\chi_{1}^{n_{1}}(\gamma_{1})+1,S_{2}-\chi_{2}^{n_{2}}(\gamma_{2})+1)$. Since we have
$(S_{1}-\chi_{1}^{n_{1}}(\gamma_{1})+1)g_{\boldsymbol{n}}$ $=(\gamma_{1}-\chi_{1}^{n_{1}}(\gamma_{1}))g_{\boldsymbol{n}}=g_{\boldsymbol{n}+(p-1,0)}\quad\text{and}$
$(S_{2}-\chi_{2}^{n_{2}}(\gamma_{2})+1)g_{\boldsymbol{n}}$ $=(\gamma_{2}-\chi_{2}^{n_{2}}(\gamma_{2}))g_{\boldsymbol{n}}=g_{\boldsymbol{n}+(0,p-1)},$
the element $f_{\boldsymbol{n}}(S_{1},S_{2})g_{\boldsymbol{n}}$ can be written as a $\mathbb{Z}_{p}[[S_{1},S_{2}]]$-linear combination of $g_{\boldsymbol{n}+(p-1,0)}$ and $g_{\boldsymbol{n}+(0,p-1)}$. By repeating this argument, it follows that $x$ can be expressed as a $\mathbb{Z}_{p}[[S_{1},S_{2}]]$-linear combination of $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}}$, where $\boldsymbol{m}\in I$ ranges over the indexes satisfying $|\boldsymbol{m}|=m+(p-1)$ and $\boldsymbol{m}\equiv\boldsymbol{m}_{0}\bmod p-1$. This concludes the proof of the claim. ∎
4.2. Group-theoretic lemmas
In this subsection, we complete a proof of Theorem 2.12. First, we prepare a series of group-theoretic lemmas. We first prove a generalization of Lemma 3.1 in [Sha02] to the case of a free pro-$p$ group of countably infinite rank:
Lemma 4.9.
Let $\mathcal{F}$ be a pro-$p$ group strongly generated by $y$ and $\{x_{i}\}_{i\geq 1}$. For each $i\geq 1$, let $x_{i,1}\coloneqq x_{i}$ and we inductively define
$x_{i,j+1}\coloneqq[y,x_{i,j}]x_{i,j}^{pa_{i,j}}$
for some $a_{i,j}\in\mathbb{Z}_{p}$ for every $j\geq 1$. Denote the normal closure of $\{x_{i}\}_{i\geq 1}$ in $\mathcal{F}$ by $H$. Then the following two assertions hold:
  1. (1)
    $H$ is strongly generated by $\{x_{i,j}\}_{i,j\geq 1}$.
  2. (2)
    If $\mathcal{F}$ is a free pro-$p$ group on $y$ and $\{x_{i}\}_{i\geq 1}$, then $H$ is a free pro-$p$ group on $\{x_{i,j}\}_{i,j\geq 1}$.
Proof.
Let $\mathcal{K}$ be a free pro-$p$ group on the set $\{\tilde{x}_{i,j}\}_{i,j\geq 1}$. We define a two-variable filtration on this group by
$\mathcal{K}_{i,j}\coloneqq\langle\tilde{x}_{i^{\prime},j^{\prime}}\mid i^{\prime}\geq i\text{ or }j^{\prime}\geq j\rangle_{\rm{normal}}$
for every $i,j\geq 1$. The quotient $\mathcal{K}/\mathcal{K}_{i,j}$ is a free pro-$p$ group of finite rank, and the image of $\{\tilde{x}_{i,j}\}_{i^{\prime}<i,j^{\prime}<j}$ forms a basis of this quotient.
[…]
Let us set $M\coloneqq\mathcal{K}^{\rm{ab}}/p$, and let $M_{i,j}\subset M$ denote the image of $\mathcal{K}_{i,j}$ inside $M$. Since taking the maximal abelian quotients and taking reductions modulo $p$ are both right exact, the quotient $M/M_{i,j}$ is naturally isomorphic to $(\mathcal{K}/\mathcal{K}_{i,j})^{\rm{ab}}/p$. Note that the kernel of the homomorphism
[…]
is pro-$p$, since the kernel of each $\mathrm{Aut}(\mathcal{K}/\mathcal{K}_{i,j})\to\mathrm{Aut}(M/M_{i,j})$ is so. Therefore, to prove the claim, it suffices to show that the homomorphism ^$\hat{\mathbb{Z}}\to\mathrm{Aut}_{\rm{fil}}(M)$ corresponding to the image of $\phi$ factors through $\mathbb{Z}_{p}$.
By the construction of the automorphism $\phi$, it follows that
$\phi^{p^{n}}\in\ker\left(\mathrm{Aut}_{\rm{fil}}(M)\to\mathrm{Aut}(M/M_{p^{n},p^{n}})\right)$
for every $n\geq 1$. In fact, a direct computation shows that $\phi^{p}(\tilde{x}_{i,j})=\tilde{x}_{i,j+p}+\tilde{x}_{i,j}$ in $M$ for every $i,j$, and iterating $\phi^{p}$ verifies the claim. Hence the image of $\phi$ in $\mathrm{Aut}_{\rm{fil}}(M)=\mathrm{Aut}(M/M_{p^{n},p^{n}})$ is a pro-$p$ group. It follows that the map $\mathbb{Z}\to\mathrm{Aut}_{\rm{fil}}(\mathcal{K})$ corresponding to $\phi$ naturally extends to a continuous homomorphism $\mathbb{Z}_{p}\to\mathrm{Aut}_{\rm{fil}}(\mathcal{K})$, and we can take the associated semi-direct product $\mathcal{K}\rtimes\mathbb{Z}_{p}$.
There is a unique homomorphism $\mathcal{K}\to H$ sending $\tilde{x}_{i,j}$ to $x_{i,j}$ for every $i,j\geq 1$ by the freeness of $H$. Since the action of $\phi$ on $\mathcal{K}$ is compatible with the conjugation by $y$ on $H$, we can extend this homomorphism to $\mathcal{K}\rtimes\mathbb{Z}_{p}\to\mathcal{F}$ by mapping $\phi$ to $y$, and the resulting homomorphism is surjective by construction. Hence $\mathcal{K}\to H$ is also surjective. This proves the assertion of (1). Moreover, if $\mathcal{F}$ is a free pro-$p$ group on $y$ and $\{x_{i}\}_{i}$, the universal property of $\mathcal{F}$ leads to the existence of the inverse of $\mathcal{K}\rtimes\mathbb{Z}_{p}\to\mathcal{F}$ constructed above. Hence $H$ is isomorphic to $\mathcal{K}$, proving the second assertion. ∎
Corollary 4.10.
Let $r\geq 1$ be an integer and $\mathcal{F}$ be a pro-$p$ group generated by $y_{1},y_{2}$ and $\{x_{i}\}_{1\leq i\leq r}$. For each $1\leq i\leq r$, let $x_{i,(0,0)}\coloneqq x_{i}$ and we inductively define
$x_{i,(j+1,0)}\coloneqq[y_{1},x_{i,(j,0)}]x_{i,(j,0)}^{pa_{i,j}}$
for some $a_{i,j}\in\mathbb{Z}_{p}$ and $j\geq 0$. Similarly, for each $1\leq i\leq r$ and each $j\geq 0$, define
$x_{i,(j,k+1)}\coloneqq[y_{2},x_{i,(j,k)}]x_{i,(j,k)}^{pb_{i,j,k}}$
for some $b_{i,j,k}\in\mathbb{Z}_{p}$ and $k\geq 0$. Moreover, let $z_{(0,0)}\coloneqq[y_{1},y_{2}]$ and define
$z_{(i+1,0)}\coloneqq[y_{1},z_{(i,0)}]z_{(i,0)}^{p\alpha_{i}}$
for some $\alpha_{i}\in\mathbb{Z}_{p}$ and $i\geq 0$. Finally, for each $i\geq 0$, we define
$z_{(i,j+1)}\coloneqq[y_{2},z_{(i,j)}]z_{(i,j)}^{p\beta_{i,j}}$
for some $\beta_{i,j}\in\mathbb{Z}_{p}$ and each $j\geq 0$. Let $H$ denote the normal closure of $\{x_{i}\}_{1\leq i\leq r}$ and $z_{(0,0)}$. Then $H$ is strongly generated by $\{x_{i,(j,k)}\}_{\begin{subarray}{c}1\leq i\leq r,\\ j,k\geq 0\end{subarray}}$ and $\{z_{i,j}\}_{i,j\geq 0}$.
Proof.
We may assume that $\mathcal{F}$ is a free pro-$p$ group on $y_{1},y_{2}$ and $\{x_{i}\}_{1\leq i\leq r}$. First, let $\mathcal{F}_{1}$ be the kernel of the following homomorphism:
$\mathcal{F}\to\mathbb{Z}_{p};\;y_{1}\mapsto 1,y_{2}\mapsto 0,x_{i}\mapsto 0\quad(1\leq i\leq r)$
Then, by Lemma 4.9, the group $\mathcal{F}_{1}$ is a free pro-$p$ group on $y_{2}$, $\{z_{(i,0)}\}_{i\geq 0}$ and $\{x_{i,(j,0)}\}_{\begin{subarray}{c}1\leq i\leq r,\\ j\geq 0\end{subarray}}$. Moreover, the group $H$ coincides the kernel of the homomorphism
$\mathcal{F}_{1}\to\mathbb{Z}_{p};\;y_{2}\mapsto 1,z_{(i,0)}\mapsto 0\;(i\geq 0)\text{ and }x_{i,(j,0)}\mapsto 0\;(1\leq i\leq r,j\geq 0).$
By applying Lemma 4.9 again, it follows that $H$ is a free pro-$p$ group on $\{z_{i,j}\}_{i\geq 0,j\geq 0}$ and $\{x_{i,(j,k)}\}_{\begin{subarray}{c}1\leq i\leq r\\ j,k\geq 0\end{subarray}}$. This concludes the proof. ∎
The following lemma is used to compare the filtration $\{F^{m}G\}_{m\geq 1}$ on $G$ with a certain canonical filtration on $G$ associated to $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$.
Lemma 4.11.
Let $\mathcal{G}$ be a free pro-$p$ group on the set $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$.
  1. (1)
    There exists a unique descending central filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ on $\mathcal{G}$ satisfying the following property: $(\mathrm{i})$ $\tilde{\sigma}_{\boldsymbol{m}}\in\tilde{F}^{\lvert\boldsymbol{m}\rvert}\mathcal{G}$ for every $\boldsymbol{m}\in I$. $(\mathrm{ii})$ If $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is a descending central filtration satisfying $(\mathrm{i})$, then $\tilde{F}^{m}\mathcal{G}\subset F^{m}\mathcal{G}$ for every $m\geq 1$.
  2. (2)
    The graded Lie algebra $\bigoplus_{m\geq 1}\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G}$ is freely generated by the image of $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$.
  3. (3)
    The intersection $\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}$ is trivial.
Proof.
We construct $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ as follows: First, let $\tilde{F}^{1}\mathcal{G}\coloneqq\mathcal{G}$. For $m\geq 2$, we inductively define $\tilde{F}^{m}\mathcal{G}$ as
$\tilde{F}^{m}\mathcal{G}\coloneqq\langle\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\lvert\boldsymbol{m}\rvert\geq m},\{[\tilde{F}^{m^{\prime}}\mathcal{G},\tilde{F}^{m^{\prime\prime}}\mathcal{G}]\}_{\begin{subarray}{c}m^{\prime}<m,m^{\prime\prime}<m\\ m\leq m^{\prime}+m^{\prime\prime}\end{subarray}}\rangle_{{\rm normal}}.$
Since $[\tilde{F}^{m}\mathcal{G},\tilde{F}^{1}\mathcal{G}]\subset\tilde{F}^{m}\mathcal{G}$, it follows that $\tilde{F}^{m+1}\mathcal{G}\subset\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Apparently, the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ defines a descending central filtration on $\mathcal{G}$ satisfying the condition ($\mathrm{i}$) in the assertion of (1). Moreover, if $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is a descending central filtration satisfying ($\mathrm{i}$), then we have $F^{1}\mathcal{G}=\tilde{F}^{1}\mathcal{G}=\mathcal{G}$. By induction on $m$, it follows that $\tilde{F}^{m}\mathcal{G}\subset F^{m}\mathcal{G}$ holds for every $m\geq 1$. Hence the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ also satisfies ($\mathrm{ii}$). The uniqueness is clear. The proof of the assertion of (2) is similar to that of [Iha02, p.263, 5].
[…]
For an integer $m\geq n$, the image of $\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ is normally generated by the image of
$\{[\tilde{F}^{m^{\prime}}\mathcal{G},\tilde{F}^{m^{\prime\prime}}\mathcal{G}]\}_{\begin{subarray}{c}m^{\prime}<m,m^{\prime\prime}<m\\ m\leq m^{\prime}+m^{\prime\prime}\end{subarray}}.$
We claim that the image of $\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ contained in the $r_{m}$-th component $F_{n}(r_{m})$ of the descending central series of $F_{n}$, where an integer $r_{m}$ is defined by
$r_{m}\coloneqq\Bigl{\lfloor}\frac{m}{n}\Bigr{\rfloor}+1$
for every $m\geq 1$. Once this claim is proved, then the image of $\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ is contained in the intersection $\cap_{m\geq n}F_{n}(r_{m})=\{1\}$, and hence the assertion follows.
The claim holds for every $m\leq 2n-1$. Assume that the claim also holds for every $m\leq kn-1$ for some $k\geq 2$ and we prove the claim for $m=kn,kn+1,\dots(k+1)n-1$ in order. Write $m=kn+r$ for some $10\leq r\leq n-$ 1.
Let $m^{\prime},m^{\prime\prime}$ be positive integers less than $m$ satisfying $m^{\prime}+m^{\prime\prime}\geq m$, and write $m^{\prime}=k^{\prime}n+r^{\prime}$, $m^{\prime}=k^{\prime\prime}n+r^{\prime\prime}$ for some $0\leq k^{\prime},k^{\prime\prime}\leq k$ and $10\leq r^{\prime},r^{\prime\prime}\leq n-$ 1. Since
$m^{\prime}+m^{\prime\prime}=(k^{\prime}+k^{\prime\prime})n+(r^{\prime}+r^{\prime\prime})\geq m=kn+r,$
It holds that
$r_{m^{\prime}}+r_{m^{\prime\prime}}=(k^{\prime}+k^{\prime\prime})+2\geq(k+1)+\frac{n+r-(r^{\prime}+r^{\prime\prime})}{n}.$
Since $\frac{n+r-(r^{\prime}+r^{\prime\prime})}{n}\geq\frac{2-n}{n}>-1$, it holds that
$r_{m^{\prime}}+r_{m^{\prime\prime}}\geq r_{m}.$
Therefore, the image of $[\tilde{F}^{m^{\prime}}\mathcal{G},\tilde{F}^{m^{\prime\prime}}\mathcal{G}]$ is contained in
$[F_{n}(r_{m^{\prime}}),F_{n}(r_{m^{\prime\prime}})]\subset F_{n}(r_{m^{\prime}}+r_{m^{\prime\prime}})\subset F_{n}(r_{m}).$
Hence the image of $\tilde{F}^{m}\mathcal{G}$ is contained in $F_{n}(r_{m})$, as desired. ∎
The next lemma gives an explicit set of generators of $G=\mathrm{Gal}(\Omega/K(p))$:
Lemma 4.12.
The group $G$ is generated by $\gamma_{1}$, $\gamma_{2}$ and $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}\setminus\{(p-1,p-1)\}}$.
Proof.
We use Lemma 4.2 and definitions of $\sigma_{\boldsymbol{m}}$. It follows that the image of $\sigma_{\boldsymbol{m}}$ in $A=\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$ generates $A^{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$. Moreover, the image of $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$ also generates $A^{(1,1)}$. Finally, recall that $\sigma_{(p-1,p-1)}$ coincides with the commutator $[\gamma_{1},\gamma_{2}]$. We have proved in the proof of Lemma 4.4 that this element generates $A^{(p-1,p-1)}$.
Hence it follows that the $\Lambda$-module $A$ is generated by the image of $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, which is equivalent to saying that $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is normally generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$ (as a normal subgroup of $G$). Hence $\gamma_{1}$, $\gamma_{2}$ and $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}\setminus\{(p-1,p-1)\}}$ generates the group $G$, as desired. ∎
Finally, we give a proof of Theorem 2.12.
Proof of Theorem 2.12.
By Lemma 4.12, the group $G$ is generated by
$\gamma_{1},\gamma_{2}\quad\text{and}\quad\{\sigma_{\boldsymbol{m}}(=g_{\boldsymbol{m}})\}_{\boldsymbol{m}\in I_{0}\setminus\{(p-1,p-1)\}}.$
Now we recall the discussion occurring at constructing $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ for $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$. There were two ways to construct them, starting from $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$, but they yield the same elements on $A=\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$. Hence, by applying Corollary 4.10 to $\mathcal{F}=G$, $y_{1}=\gamma_{1}$, $y_{2}=\gamma_{2}$ and $\{x_{i}\}_{1\leq i\leq r}=\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, it follows that $F^{1}G=\mathrm{Gal}(\Omega/K(p^{\infty}))$ is strongly generated by $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$. By Lemma 4.7, it follows that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ also strongly generates $F^{1}G$.
We claim that the projection $F^{1}G\to F^{1}G^{\ast}$ is an isomorphism, which is equivalent to the desired equality $\Omega=\Omega^{\ast}$. Let $\mathcal{G}$ be a free pro-$p$ group on the set $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, $\varphi\colon\mathcal{G}\to F^{1}G$ a surjective homomorphism sending $\tilde{\sigma}_{\boldsymbol{m}}$ to $\sigma_{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I$ and $\varphi^{\ast}\colon\mathcal{G}\to F^{1}G^{\ast}$ the composite of $\varphi$ with the projection $F^{1}G\to F^{1}G^{\ast}$.
The group $\mathcal{G}$ comes equipped with a descending central filtration $\{F^{m}\mathcal{G}\}_{m\geq 1}$ induced by $\{F^{m}G^{\ast}\}_{m\geq 1}$ through the map $\varphi^{\ast}$. Note that the graded Lie algebra associated to $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is naturally isomorphic to $\mathfrak{g}$, and we have
$\ker(\varphi^{\ast})=(\varphi^{\ast})^{-1}(\cap_{m\geq 1}F^{m}G^{\ast})=\cap_{m\geq 1}F^{m}\mathcal{G}.$
Hence, to show the injectivity of the map $\varphi^{\ast}$, it suffices to prove that the intersection of the induced filtration on $\mathcal{G}$ is trivial. By Lemma 4.11 (1), the group $F^{m}\mathcal{G}$ contains $\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Hence we have the following commutative diagram:
${\left(\bigoplus_{m\geq 1}\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G}\right)\otimes\mathbb{Q}_{p}}$${\left(\bigoplus_{m\geq 1}F^{m}\mathcal{G}/F^{m+1}\mathcal{G}\right)\otimes\mathbb{Q}_{p}\cong\mathfrak{g}\otimes\mathbb{Q}_{p}}$${\bigoplus_{m\geq 1}\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G}}$${\bigoplus_{m\geq 1}F^{m}\mathcal{G}/F^{m+1}\mathcal{G}.}$
Since we assume the analogue of the Deligne-Ihara conjecture (Conjecture 2.8), it holds that $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generates $\bigoplus_{m\geq 1}(F^{m}\mathcal{G}/F^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ as a graded Lie algebra. However, by Lemma 4.11 (2), the Lie algebra $\bigoplus_{m\geq 1}(\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ is also generated by $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which forces the upper horizontal arrow in the diagram to be an isomorphism. As a consequence, it holds that the lower horizontal arrow in the diagram is injective. As is observed by induction on $m$, this injectivity is equivalent to saying that two filtrations $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ and $\{F^{m}\mathcal{G}\}_{m\geq 1}$ coincide with each other. Hence by Lemma 4.11 (3), we have
$\cap_{m\geq 1}F^{m}\mathcal{G}=\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}=\{1\},$
showing the injectivity of the map $\varphi^{\ast}\colon\mathcal{G}\to F^{1}G^{\ast}$. Hence the map $\varphi^{\ast}$ is an isomorphism, so is the natural surjection $F^{1}G\to F^{1}G^{\ast}$, as desired. ∎
Remark 4.13.
(1) As a byproduct of the proof, the elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generate the graded Lie algebra $\mathfrak{g}$ over $\mathbb{Z}_{p}$, by Lemma 4.11 (2). In other words, under the assumption of Theorem 2.12, the integral version of Conjecture 2.8 holds. In the same way that the integral version of the Deligne-Ihara conjecture is related to generalized Greenberg’s conjecture for $\mathbb{Q}(\mu_{p})$ [Sha02, Theorem 1.3], the integral version of Conjecture 2.8 is related to generalized Greenberg’s conjecture for $K(p)$. We hope to discuss this topic in a future paper.
(2) As another byproduct, the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is found to be a free pro-$p$ group generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$. This consequence is consistent with Remark 4.3 (2).
Appendix A
In this appendix, we keep the same notation as in the main text, i.e. let $K$ be an imaginary quadratic field of class number one, $p$ a prime $\geq 5$ which splits in $K$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ and $\mu_{p}$ the group of $p$-th roots of unity. Moreover, let $K_{\infty}$ be the unique $\mathbb{Z}_{p}^{2}$-extension of $K$. In this appendix, we determine the structure of the Galois group of the maximal pro-$p$ extension $\Omega^{\mathrm{cyc}}_{K}$ of $K(\mu_{p})$ unramified outside $p$ when $p$ satisfies a certain assumption.
We denote a unique prime of $K(\mu_{p})$ lying above $\mathfrak{p}$ (resp. $\bar{\mathfrak{p}}$) by the same letter, and fix an arbitrary prime of $\Omega_{K}^{\mathrm{cyc}}$ above $\mathfrak{p}$. Then we have the homomorphism
$\phi_{\mathfrak{p}}:G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)}\to\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p})),$
associated to the decomposition group of the fixed prime of $\Omega_{K}^{\mathrm{cyc}}$.
The Galois group $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is said to be local type if $\phi_{\mathfrak{p}}$ is surjective [NSW08, (10.9.6) Definition]. Moreover, it is called purely local if $\phi_{\mathfrak{p}}$ is an isomorphism [Win90]. The aim of this appendix is to prove the following theorem, which tells us exactly when $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is purely local:
Theorem A.1.
The Galois group $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is purely local if and only if $p$ does not divide the class number of $K(\mu_{p})$ and there exists a unique prime of $K_{\infty}$ lying above $\mathfrak{p}$.
Remark A.2.
(1) The Galois group $G_{\mathbb{Q}(\mu_{p})}^{(p)}$ is a pro-$p$ Demushkin group of rank $p+1$ [NSW08, (7.5.11) Theorem (ii)]. In particular, it is generated by $p+1$ elements which are subject to a single relation. On the other hand, if $p$ does not divide the class number of $K(\mu_{p})$, then $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is also generated by $p+1$ elements satisfying a single relation [NSW08, (10.7.13) Theorem]. Hence the condition on the class number implies that the concerned Galois groups have the same numbers of generators and relations.
(2) The latter condition on the number of primes of $K_{\infty}$ above $\mathfrak{p}$ fails when $K=\mathbb{Q}(\sqrt{-1})$ and $p=29789$ [Ish25, Example 4.16]. This is the only example the author has been able to find where the latter condition does not hold when $K=\mathbb{Q}(\sqrt{-1})$.
Proof of Theorem A.1.
Assume that $\phi_{\mathfrak{p}}$ is an isomorphism. Then there exists a unique prime of $\Omega^{\mathrm{cyc}}_{K}$ above $\mathfrak{p}$. Since we have $K_{\infty}\subset K_{\infty}K(\mu_{p})\subset\Omega^{\mathrm{cyc}}_{K}$, a prime of $K_{\infty}$ lying above $\mathfrak{p}$ is also unique. Moreover, the class number of $K(\mu_{p})$ is not divisible by $p$ by [Ish25, Proposition 4.13].
Conversely, assume that $p$ does not divide the class number of $K(\mu_{p})$ and there exists a unique prime of $K_{\infty}$ above $\mathfrak{p}$. By [Ish25, Proposition 4.13], it suffices to show that $\phi_{\mathfrak{p}}$ is surjective. As a first step, we consider the kernel of the restriction map
$H^{1}(\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p})\to H^{1}(G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)},\mu_{p})\xrightarrow{\sim}\mathbb{Q}_{p}(\mu_{p})^{\times}/p,$
induced by $\phi_{\mathfrak{p}}$. Since the class number of $K(\mu_{p})$ is prime to $p$, we have
$O_{K(\mu_{p})}[1/p]^{\times}/p\xrightarrow{\sim}H^{1}(\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p})$
by using a similar argument as the proof of Lemma 4.2. Hence the surjectivity of $\phi_{\mathfrak{p}}$ is equivalent to the injectivity of the map
$\iota_{\mathfrak{p}}:O_{K(\mu_{p})}[1/p]^{\times}/p\to\mathbb{Q}_{p}(\mu_{p})^{\times}/p$
induced by the $\mathfrak{p}$-adic completion. The map $\iota_{\mathfrak{p}}$ is compatible with the action of $\mathrm{Gal}(K(\mu_{p})/K)$, the decomposition group of $\mathfrak{p}$. Moreover, if we define
$\iota_{\bar{\mathfrak{p}}}:O_{K(\mu_{p})}[1/p]^{\times}/p\to\mathbb{Q}_{p}(\mu_{p})^{\times}/p$
to be the map induced by the $\bar{\mathfrak{p}}$-adic completion, then the map
$(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}}):O_{K(\mu_{p})}[1/p]^{\times}/p\to\mathbb{Q}_{p}(\mu_{p})^{\times}/p\times\mathbb{Q}_{p}(\mu_{p})^{\times}/p.$
is injective, since the dimension of the kernel of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ is equal to that of the $p$-class group of $K(\mu_{p})$ modulo $p$ (see [NSW08, Proof of (10.7.3)], for example).
The injectivity of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ already implies that the kernel of $\iota_{\mathfrak{p}}$ is at most a single Tate twist $\mathbb{F}_{p}(1)$: More precisely, we show that the composite
$\iota_{\mathfrak{p}^{+}}:O_{K(\mu_{p})^{+}}[1/p]^{\times}/p\to O_{K(\mu_{p})}[1/p]^{\times}/p\xrightarrow{\iota_{\mathfrak{p}}}\mathbb{Q}_{p}(\mu_{p})^{\times}/p.$
is injective, where $K(\mu_{p})^{+}$ denotes the maximal real subfield of $K(\mu_{p})$. Note that
$\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})^{+}}[1/p]^{\times}/p=p\quad\text{and}\quad\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})}[1/p]^{\times}/p=p+1$
by Dirichlet’s unit theorem. Moreover, the cokernel of the inclusion
$O_{K(\mu_{p})^{+}}[1/p]^{\times}/p\to O_{K(\mu_{p})}[1/p]^{\times}/p$
is generated by the image of $\mu_{p}$. The injectivity of $\iota_{\mathfrak{p}^{+}}$ follows from the that of the map $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ and the fact that a prime of $K(\mu_{p})^{+}$ lying above $p$ is unique, since the restriction of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ on $O_{K(\mu_{p})^{+}}[1/p]^{\times}/p$ is just the diagonal map which is induced by the completion at the unique prime $K(\mu_{p})^{+}$ lying above $p$.
Now it is easy to check by using [NSW08, (8.7.2) Proposition] that
$O_{K(\mu_{p})}[1/p]^{\times}/p\cong\mathbb{F}_{p}^{\oplus 2}\oplus\mathbb{F}_{p}(1)^{\oplus 2}\oplus\bigoplus_{m=2}^{p-2}\mathbb{F}_{p}(m)$
as an $\mathbb{F}_{p}[\mathrm{Gal}(K(\mu_{p})/K)]$-module. Since $\ker(\iota_{\mathfrak{p}})$ is stable under the action of $\mathrm{Gal}(K(\mu_{p})/K)$, the injectivity of $\iota_{\mathfrak{p}^{+}}$ implies that $\ker(\iota_{\mathfrak{p}})\subset\mathbb{F}_{p}(1)^{\oplus 2}$. In terms of Galois groups, the cokernel of the map
$G_{\mathbb{Q}_{p}(\mu_{p})}^{(p),\mathrm{ab}}/p\to\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))^{\mathrm{ab}}/p$
induced by $\phi_{\mathfrak{p}}$ is contained in $\mathbb{F}_{p}^{2}$ (with trivial $\mathrm{Gal}(K(\mu_{p})/K)$-action). However, since the $\mathbb{F}_{p}^{2}$-extension of $K(\mu_{p})$ corresponding to $\mathbb{F}_{p}(1)^{\oplus 2}$ is given by the quotient
$\mathrm{Gal}(K_{\infty}K(\mu_{p})/K(\mu_{p}))\cong\mathbb{Z}_{p}^{2}\to\mathbb{F}_{p}^{2}$
and we have assumed that there is a unique prime of $K_{\infty}K(\mu_{p})$ lying above $\mathfrak{p}$, the decomposition group at $\mathfrak{p}$ must coincide with $\mathbb{F}_{p}^{2}$. Hence $\phi_{\mathfrak{p}}$ is surjective as desired. ∎
Theorem A.1 allows us to use the local Tate duality to compute cohomology groups of the global Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p}))$. Consequently, we obtain the following corollary:
Corollary A.3.
Under the same assumption as in Theorem A.1, the cohomology group $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(m_{1},m_{2}))$ is finite for every $(m_{1},m_{2})\in I$ such that $m_{1}\equiv m_{2}\bmod p-1$. Consequently, the $(m_{1},m_{2})$-th elliptic Soulé character $\kappa_{(m_{1},m_{2})}$ is nontrivial for every such $(m_{1},m_{2})\in I$.
Proof.
The assertion follows from Theorem A.1 and Theorem 2.5. ∎